### v1

# GIM: A Million-scale Benchmark for Generative Image Manipulation Detection and Localization

Abstract
The extraordinary ability of generative models emerges as a new trend in image editing and generating realistic images, posing a serious threat to the trustworthiness of multimedia data and driving the research of image manipulation detection and location(IMDL). However, the lack of a large-scale data foundation makes IMDL task unattainable. In this paper, a local manipulation pipeline is designed, incorporating the powerful SAM, ChatGPT and generative models. Upon this basis, We propose the GIM dataset, which has the following advantages: 1) Large scale, including over one million pairs of AI-manipulated images and real images. 2) Rich Image Content, encompassing a broad range of image classes 3) Diverse Generative Manipulation, manipulated images with state-of-the-art generators and various manipulation tasks. The aforementioned advantages allow for a more comprehensive evaluation of IMDL methods, extending their applicability to diverse images. We introduce two benchmark settings to evaluate the generalization capability and comprehensive performance of baseline methods. In addition, we propose a novel IMDL framework, termed GIMFormer, which consists of a ShadowTracer, Frequency-Spatial Block (FSB), and a Multi-window Anomalous Modelling (MWAM) Module. Extensive experiments on the GIM demonstrate that GIMFormer surpasses previous state-of-the-art works significantly on two different benchmarks. Code page: GIM.
Keywords:
Dataset and Benchmark Image Manipulation Detection and Location Forgery Traces Extraction Multi-scale Feature Learning
11footnotetext: * These authors contributed equally to this research, which was done during Yirui Chen and Quan Zhang’s internship at Huawei Noah’s Ark Lab. ‡ Project lead † Corresponding authors
1 Introduction
Figure 1: Example image from the GIM dataset. Our dataset includes images manipulated by three state-of-the-art generators: Stable-Diffusion, GLIDE, and DDNM. The first column shows authentic images, while the second column presents forgery masks. The third column displays forged images.
Images are one of the most essential media for information transmission in modern society, which is widely spread on public platforms such as news and social media. With the rapid advancement of generative methods [11, 52, 16, 4], such natural information can be more easily tampered with for specific purposes such as tampering with an object or person. The image of this class is particularly convincing since its visual comprehensibility, leads to serious information security risks in social areas. For instance, an AI-generated image of black smoke billowing from what appeared to be a government building near the Pentagon set off investor fears, sending stocks tumbling [46]. Therefore, it is of utmost urgency to develop effective methods to examine whether an image is modified by generative models, meanwhile identifying the exact location of the tampering. However, the traditional image manipulation detection and location(IMDL) dataset [55, 13, 60] overlook the powerful generative models and generative IMDL dataset are limited with scale, thus not sufficient to comprehensively evaluate the performance of IMDL methods and benefits the IMDL community.
To this end, we propose a million-level generative-based IMDL dataset, termed GIM, to provide a reliable database for AI Generated Content (AIGC) security. GIM leverage the diffusion model [44, 52, 59] and SAM [32, 51], with ImageNet [10] and VOC [14] as the input. SAM is leveraged to locate the tampering region, and diffusion models paint the reasonable content in the tampering area. The excellent diffusion model ensures dataset fidelity and rich categories guarantee the diversity of images in GIM. GIM contains a total of over one million generative tampered images. Based on this, the IMDL methods are evaluated and benchmarked. To develop an appropriate benchmark scale, we explore the impact of different amounts of training manipulated data. The final benchmark contains about 300k manipulated images with their tampering masks. To simulate the real situation, we investigate the effect of degradation and apply the images to the random degradation method. Two benchmark settings are designed to evaluate the comprehensive and generalization. In addition, we recreated existing IMDL methods in a fair manner, providing the basis for future development. Overall, as a generative IMDL benchmark, GIM possesses the following advantages: 1) GIM has a large and reasonable data scale, including rich image categories and contents. 2) GIM contains various generative manipulation models and tasks. 3) GIM proposes two settings for verifying generalization capability and comprehensive performance in existing methods.
The existing methods emphasize traditional image manipulations also known as “cheapfakes”. Generative manipulations introduce lethal alterations in content with no apparent frequency or structural inconsistency. To address the above issue, we introduce GIMFormer, a transformer-based framework for generative IMDL. The ShadowTracer is designed to embed the nuanced artifacts inherent in generative tampering and serves as prior information. The Frequency-Spatial Block harvests the manipulation clues in frequency and spatial domains. Furthermore, the Multi Windowed Anomalous Modelling Module captures local inconsistencies at different scales to refine the features. By doing so, our model extracts features from both the RGB and learned tampering traces map, captures details from the frequency domain and spatial domain, and models inconsistencies at different scales for precise manipulation detection and location.
We conduct experiments on our proposed GIM. Both the qualitative and quantitative results demonstrate that GIMFormer can significantly outperform previous state-of-the-art methods in terms of both comprehensive capability and generalization ability.
In summary, our main contributions are as follows:
  • We build a local manipulation pipeline and construct a comprehensive large-scale dataset with various SOTA generative models and manipulation tasks to further facilitate the research on the IMDL task.
  • We investigate the impact of data scales and degradation on IMDL tasks to select an appropriate scale and simulate real-world scenarios. Based on this, we construct a reasonable benchmark and two settings to evaluate the generalization and comprehensive capabilities of IMDL methods.
  • We propose GIMFormer, a generative IMDL framework consisting of ShadowTracer to embed subtle traces, Frequency-Spatial Block to extract forgery features in the frequency and spatial domains and Multi-Windowed Anomalous Modelling Module to capture pixel-level inconsistencies at multi scales.
  • Extensive experiments show that GIMFormer achieves state-of-the-art IMDL performance in terms of generalization and comprehensive capabilities.
2 Related Work
2.1 Image Forensic Datasets
The research community has dedicated significant efforts to establishing robust datasets for image forensics. Early datasets primarily focus on one type of manipulation. The Columbia dataset  [43, 24] is the pioneering dataset focusing on splicing forgeries. CASIA v1.0 and CASIA v2.0  [13] are the first to incorporate multiple types of manipulations within a single dataset, with forged images manually crafted using Adobe Photoshop. The Wild Web [66] dataset collects forged images from the Internet, significantly surpassing the scale of previous datasets. NIST [17] Datasets provide an extensive collection of datasets serving as a crucial standard for assessing media tampering detection methods. DEFACTO [42]is automatically generated from the Microsoft COCO dataset, encompassing four categories of forgeries: Copy and Move, Splicing, Object Removal, and Morphing. IMD2020 [45] provides a range of locally manipulated images generated through manual operations or random slicing, while also collecting online images without apparent manipulation traces. Recently, HiFi-IFDL [20] construct a hierarchical fine-grained dataset containing some representative forgery methods. AutoSplice [29] leverage the capabilities of large-scale language-image models like DALL-E2  [48] to facilitate automatic image editing and generation. CocoGlide [19] contains 512 images using the GLIDE diffusion model. Some datasets focused exclusively on facial manipulations [54, 18, 9] or entirely synthesized images[71, 56, 3].
However, these datasets have certain limitations, such as small data sizes and limited manipulation techniques or tasks. Recent advances in generative models [44, 23, 52] have demonstrated remarkable abilities in generative-based manipulation. These developments have led to the emergence of open-source projects [51, 65, 15] for manipulation. Leveraging the power of these models, we introduce GIM, a large-scale generative-based manipulation dataset.
Figure 2: An overview of the dataset generation. With the input of the original image and the desired region of interest, classification attribution or mouse input as a User query, the manipulation mask is extracted by the zero-shot segmentation model SAM. Besides, interacting with ChatGPT, we organize the tampering prompt by combining replacement classes. The final generations are produced by generative models with the input of the original image, tampering mask and tampering prompts.
2.2 Image Manipulation Detection and Localization
Early studies on natural image manipulation localization mainly focus on detecting a specific type of manipulation [7, 27, 34, 50]. Due to the exact manipulation type is unknown in real-world scenarios, most state-of-the-art methods [40, 19, 58, 64, 38, 28, 61] primarily concentrate on general manipulation. RGB-N [70] propose a dual-stream network for manipulation detection. One stream focuses on extracting RGB features and the other leverages noise features to model the disparities. MVSS-Net [12] utilizes a two-stream CNN to extract noise features and employs the Dual Attention Module to merge the output of the two-stream CNN. PSCC-Net  [40] extracts hierarchical features with a top-down path and detects whether the input image has been manipulated using a bottom-up path. ObjectFormer [58] uses object prototypes to model object-level consistencies and find patch-level inconsistencies to detect the manipulation. However, these methods focus on "cheapfake" detection and encounter challenges when applied to generative manipulation due to the tampering traces are subtle and the lack of an understanding of generative manipulation patterns. Certain works [26, 6, 9, 39, 2] focus on localizing manipulations using generative models, yet these methods primarily concentrate on human faces.
In this work, we leverage a deep network to embed the subtle artifacts inherent in generative tampering as a prior trace map. we then employ a dual network to fuse the trace map and RGB image, combining frequency and spatial information and capturing pixel-level inconsistencies at multiple scales to detect and locate the manipulation.
Table 1: Summary of previous image manipulation datasets and GIM. We showcase the number of data entries within each dataset, the image content, image sizes and the manipulation techniques they encompass. $\star$ denotes that HiFi-IFDL is composed of multiple datasets, including entirely synthesized images, traditionally manipulated images and fake face images.
Dataset Image Content Image Size Num. Images Manipulations Category
# Authetic Image # Forged Image Traditional Gen.Reprint Gen.Removal
FaceForensics++ [54] Face 480p-1080p 1000 4000
DFDC [9] Face 240p-2160p 19,154 100,000
DeeperForensics [30] Face 1080p 50,000 10,000
Columbia Gray [43] General 128 $\times$ 128 933 912
CASIA V2.0  [13] General 320 $\times$ 240-800 $\times$ 600 7,200 5,123
IMD2020 [45] General 1062 $\times$ 866 35,000 35,000
Coverage [60] General 400 $\times$ 486 100 100
AutoSplice  [29] General 256 $\times$ 256 - 4232 $\times$ 4232 3,621 2,273
HiFi-IFDL[20] General - - 1,000,000$\star$
CocoGlide [19] General 256$\times$256 - 512
GIM General 64$\times$64-6000$\times$3904 1,140,000 1,140,000
3 Dataset and Benchmark Construction
In Section 3.1, we propose an automatic data synthesis pipeline to produce generative manipulated images efficiently with unlabeled images. Leveraging this pipeline, we construct a comprehensive large-scale dataset. To build a reasonable scale benchmark for evaluating the IMDL methods, we delve into the impact of training data scale in Section 3.2. To emulate real-world conditions and establish an objective benchmark, the manipulated images and original authentic images are subjected to three random degradations (JPEG compression, downsampling, and Gaussian blur), as detailed in Section 3.3). GIM benchmark comprises over 320k manipulated images paired with authentic images (100 distinct labels in ImageNet for each generative model) for algorithm evaluation. The generated images are shown in Figure 1 with their original images and tampering masks. The specific composition of GIM can be found in the Supplementary Materials. In Section 3.4, we introduce the criteria and settings used to evaluate the performance of IMDL methods.
3.1 Generation and Details of the dataset
To construct a generative-based IMDL dataset, a large-scale natural dataset is urgently desired. ImageNet [10] and VOC [14] are awesome datasets, chosen as the starting point for our research due to the following advantages: 1) Large and diverse. 2) Wide category coverage. To sum up, we regard these datasets as the database for generative manipulation. In addition to possessing various scenarios, the proposed dataset utilizes a wealth of tampering methods. Specifically, we reprint the target class by generative inpainting [52, 44] method or remove the destination region using the model [59].Benefiting from open-source projects [33, 51], we develop our data generation pipeline.
Figure 2 illustrates the overall process. Firstly, with the classification attribute or User query, the local manipulation mask is extracted by the zero-shot segmentation network [33]. For reprint tampering, the image category is embedded in the replacement prompt and interacts with ChatGPT, which returns an approximate category. The approximate category is then embedded into the inpainting prompt. Combined with the original image, manipulation mask, and inpainting prompt, generative models generate the reprint generative tampered result. For removal tampering, only the original image and manipulation mask are required for the generated model. The GIM Dataset utilizes the entire ImageNet dataset and the Stable-Diffusion generator to construct a comprehensive dataset, providing a reliable database and laying the foundation for the research below.
3.2 Analysis of Benchmark Scale
With the data generation pipeline described above, we are qualified to generate millions, tens of millions, or even billions of data. Nonetheless, blindly increasing the amount of data does not improve the algorithm performance, but may lead to data redundancy. Therefore, taking data generated by the Stable Diffusion as an example, we explore the influence of data volume and category volume on baseline classification [21] and segmentation algorithms [63]. Training sets are the different scales of the training subset, and the validation is the same test set. As shown in Table 2, the metrics of classification and segmentation are improved as the dataset scale increases. When the scale reaches 180K, the algorithm performance is almost saturated. No matter whether the image class or data scale is added, the metrics are stagnant. Experiments demonstrate that increasing the amount of data or category brings negligible benefits when the data tends to be saturated. According to the analysis above, for the training data, the GIM benchmark selects 100 different labels from ImageNet for each generation model to create tampered images. For the test data, the GIM benchmark uses the entire test dataset for each generation model.
Table 2: Results of classification and segmentation on different scale subsets. Classification uses Resnet-50 and segmentation uses SegFormer-b0.
Total Num. of Image Image Classes Image per Class Metrics
[…]
2,800 10 280 0.6306 0.2550 0.7950
28,000 100 280 0.7480 0.3247 0.7995
180,000 100 1800 0.9129 0.5289 0.8699
360,000 200 1800 0.9131 0.5291 0.8701
500,000 500 1000 0.9131 0.5291 0.8703
Table 3: Influence of different degradation method on classification (Resnet-50) and segmentation (SegFormer-b0) .
Degradation Metrics
Cls.Acc Seg.F1
- 91.29 52.89
JPEG 83.10 45.74
Gaussian Blur 90.94 40.14
Downsample 88.34 37.12
3.3 Post-processing of Degradation Method
After being spread on the Internet, images will encounter various post-processings, such as JPEG compression, downsampling, Gaussian blur, etc. These transformations can pose challenges for image forensics methods. Figure 3.2 investigates the impact of degradation. The baseline models are trained on the clean data generated by the Stable-Diffusion training set and tested on the same test set with a specific degradation method. Experimental results denote that these degradation methods make identification more difficult. To shorten the distance from reality and build an objective benchmark, the random degradation method (JPEG compression, downsampling, Gaussian blur) is performed on the dataset. The details of the parameters can be found in Supplementary Materials.
3.4 Benchmark Settings
Dataset: The datasets used for the research of generative IMDL tasks are lacking. This makes direct comparisons and accurate measurement difficult. We use three cross-generator datasets GIM-SD (data from ImageNet manipulation by Stable Diffusion, similarly hereinafter), GIM-GLIDE, GIM-DDNM and a cross-distribution dataset GIM-VOC(data from VOC manipulation by Stable Diffusion) to present benchmarking results.
Metrics: We evaluate the performance of the proposed method on both the image manipulation detection task and localization task. For classification results, we use Accuracy (Cls.acc) as our evaluation metric, while for localization, the pixel-level AUC and F1 score on manipulation masks are adopted. Since binary masks and detection scores are required to compute F1 scores, we adopt the Equal Error Rate (EER) threshold to binarize them.
Settings: Two settings are proposed to verify the comprehensive and generalization performance of the algorithms. In the cross-generator generalization setting, the models are trained on the GIM-SD training set and tested on the GIM-GLIDE, GIM-DDNM and GIM-VOC test set to explore the generalization IMDL performance, as shown in Table 5. In the mix-generator comprehensive setting, the models are jointly trained on the GIM-SD, GIM-GLIDE and GIM-DDNM training set and tested on the correspondence test dataset respectively to evaluate the comprehensive IMDL performance, as shown in Table 4.
4 Method
[…]
Figure 3: GIMFormer architecture. The ShadowTracer takes the RGB image $x$ to extract the learned manipulation trace map $t$. The encoder uses both the RGB input and trace map $t$ to extract pyramidal features $F_{i}$ ($i\in[1,4]$) across four stages. The four-stage fused features $F_{i}$ are forwarded to the decoder for generative manipulation detection and location.
4.1 ShadowTracer
Prior manipulation detection methods mainly focus on ”cheapfake” scenes and rely on visible traces. These artifacts include distortions and sudden changes caused by manipulation of the image structure. However, generative tampering makes significant alterations to the content with no apparent frequency or structural inconsistency. As shown in Figure 4, these subtle traces are displayed with inherent patterns, not visible traces with inconsistent edges.
ShadowTracer aims to capture the inherent characteristics and subtle traces of the generative models. For a doctored image, our objective is to learn a mapping $g_{\phi}$ to map the tampered image to its latent disturbed pixel values, where $g_{\phi}$ represents a neural network with trainable parameters $\phi$. Our key observation is that the differences introduced by generative models in data distribution exhibit inherent patterns, and deep neural networks can attempt to reconstruct these variations. At the training stage, we generate pairs of the input image $x_{i}$ and the generative tampered image $G(x_{i})$, the manipulation trace can be calculated by $t_{i}=G(x_{i})-x_{i}$. The objective function for training $g_{\phi}$ can be formulated as
$\min_{\bm{\phi}}\Big{\{}\mathcal{L}_{r}(g_{\bm{\phi}}\left(G\left(\mathbf{x}_{i}\right))\right),t_{i})\Big{\}}$ (1)
where $\mathcal{L}_{r}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$.
Furthermore, the mapping network should satisfy two properties: detecting subtle tampering traces and being robust to various real-world image degradations. For this reason, image pairs are generated by mixing original and manipulated images and incorporating diverse degradation operations at the training stage. Specifically, given an input image $I$, we segment a portion and perform generative manipulation to obtain $I_{m}$. The alpha blendering [47] is utilized to the original and manipulated images to obscure obvious tampering traces. Following this, we subject the images to degradation operations such as blurring, and JPEG compression to obtain the final manipulated image. The network is trained on 64$\times$64 pixel patches randomly sampled from the dataset, adapting the loss in eq. 1.
4.2 Frequency-Spatial Block
When degradation operations are applied, artifacts in manipulated images are tricky to perceive. To improve the local expressive ability and effectively harvest discriminative cues in manipulated images, we design a Frequency-Spatial Bloc(FSB) to extract forgery features in the frequency and spatial domains.
Inspired by the recent work [58, 49, 37, 67], as shown in Figure 3, FSB consists of two branches: a frequency branch and a spatial branch. In the frequency branch, the input $X$ is converted into the frequency domain $\mathcal{F}_{T}(X)$ using the 2D FFT. A learnable filter $G_{i}$ is multiplied to modulate the spectrum and capture the important frequency information. Subsequently, the inverse FFT is applied to convert the feature back to the spatial domain, resulting in the extraction of frequency-aware features $X_{\text{f}}$. In the spatial branch, the input $X$ is processed through convolutional layers and activated using the LeakyReLU function in a separate network to enhance the expressiveness of the features and obtain refined spatial features $X_{\text{s}}$. Then $X_{\text{f}}$ and $X_{\text{s}}$ are concatenated and passed through convolutional layers and the LeakyReLU activation function to obtain enhanced information, which is then combined with the original input $X$ through element-wise summation.
The total process can be formulated by:
$X_{\text{f}}=\hat{\mathcal{F}}_{T}(\mathcal{F}_{T}(X)\odot G_{i})$ (2)
$X_{\text{s}}=\mathrm{Conv_{L}}\left(\mathrm{Conv}\left(X\right)\right)$
$X_{out}=\mathrm{Conv_{L}}([X_{\mathrm{f}},X_{\mathrm{s}}])+X,$
where $\odot$ denotes the Hadamard product , $\mathrm{Conv_{L}}$ denotes convolution with LeakyReLU and $\begin{bmatrix}\cdot\end{bmatrix}$ denotes a concatenation operation.
Figure 4: Subtle traces are left by generative manipulation devoid of apparent frequency or structural inconsistency. It fatally modifies the content and this pixel-level perturbation corrupts the image. ShadowTracer captures the intrinsic pattern of the generative tampering and reconstructs the underlying tampering perturbations as priori traces.
4.3 Multi Windowed Anomalous Modelling Module
Image manipulation causes discrepancies at the pixel level. Genuine pixels are expected to exhibit consistency with neighboring pixels, while manipulated pixels may deviate and display anomalies. Former works [62, 35, 28], explore modeling such local inconsistencies. To effectively capture the pixel-level inconsistency between the manipulated and real region, we introduce the Multi Windowed Anomalous Modelling module (MWAM) to model these differences at multiple scales for fine-grained features.
As shown in Figure 3, given the input feature ${F\in H\times W\times C}$, we calculate the difference between each pixel and surrounding pixels within a local window in two branches by Eq.3.
$D^{k}_{u}[i,j]=(F[i,j]-F^{k}_{u}[i,j])/\sigma^{*},$ (3)
$\sigma^{*}=\text{maximum}(\sigma(F),1e^{-5}+w_{\sigma})$
where $u\in\{a,m\}$ denotes average or maximum branches, $\sigma(F)$ is the standard deviation of F, and $w_{\sigma}$ is a learnable non-negative weight vector of the same length as $\sigma$, $F_{a}^{k}$ and $F_{m}^{k}$ are calculated from the average and maximum values of the $k\times k$ windows in each pixel.
Different sizes $k$ are selected to model the inconsistency at different scales. Then, the obtained $N=3$ different-scale $D^{k}_{a}$ and $D^{k}_{m}$ are concatenated and fed into a convolutional network to obtain an anomaly map $M_{a}$ and $M_{m}$ of the same size as the original input.
Additionally, the anomaly score mask $\hat{\bm{S_{u}}}\in H\times W$ of the feature is computed using Eq.4.
$\hat{\bm{f}_{u}}$ $=\mathrm{DW-Conv}\left(\bm{f}\right),$ (4)
$\hat{\bm{S}_{u}}$ $=\operatorname{Sigmoid}\left(\mathrm{Conv}(C,1)\left(\hat{\bm{f}_{u}}\right)\right),$
where the $\mathrm{DW}{-\operatorname{Conv}}$ means a $3\times 3$ Depth-Wise convolution layer.
The element-wise multiplication between anomaly score $\hat{\bm{S}_{u}}$ and the anomaly map $M_{u}$ capture the anomaly information. Next, we calculate the element-wise summation between the resulting anomaly-aware map and the input feature map $X$ to obtain an anomaly-sensitive feature map. The whole process can be described as:
$\hat{X}=X+\hat{\bm{S}_{a}}\times M_{a}+\hat{\bm{S}_{m}}\times M_{m}$ (5)
Table 4: Benchmarking IMDL models for manipulation detection and localization for varying mix-generator comprehensive capability. The models are trained on the whole GIM training set and test on the correspondence test dataset respectively. The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported. Params and GFLOPS denote the models’ parameters and computational complexity with Params measured in millions (M). $\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
Method Params GFLOPS GIM-SD GIM-GLIDE GIM-DDNM
Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC
ManTranet[62] 4.0 1009.7 61.08 37.48 80.83 70.99 49.11 83.29 53.99 33.12 74.94
MVSS-Net[12] 146.9 160.0 56.12 23.17 72.03 61.29 33.12 74.94 49.15 14.11 70.09
SPAN[25] 15.4 30.9 53.15 35.62 79.28 60.01 39.46 81.21 59.15 32.1 73.55
PSCC-Net[40] 4.1 107.3 52.28 31.48 83.85 66.52 53.68 86.37 56.33 41.77 85.8
[…]
SegFormer[63] 27.5 41.3 64.25 46.17 83.26 78.11 56.77 88.73 69.29 40.19 84.65
GIMFormer (Ours) 95.9 96.2 70.92 58.61 88.25 83.89 77.31 95.42 76.72 56.25 88.31
4.4 Loss Function
For manipulation detection, we adopt a light-weight backbone in  [57] on the fourth stage feature to calculate the final binary prediction $\hat{y}$. For manipulation localization, we utilize the MLP decoder  [63] as the segmentation head to obtain a predicted mask $\hat{M}$. Given the ground-truth label y and mask M, we train GIMFormer with the following objective function:
${\mathcal{L}}={\mathcal{L}}_{cls}(y,{\hat{y}})+\lambda{\mathcal{L}}_{seg}(M,{\hat{M}}),$ (6)
where both ${\mathcal{L}}_{cls}$ and ${\mathcal{L}}_{seg}$ are binary cross-entropy loss, and $\lambda=1$ is a balancing hyperparameter.
4.5 Implementation Details.
Our approach includes two separate training steps. First, we train the ShadowTracer using a synthesized dataset of ImageNet. This training process follows a similar data generation method as described in Section 4.1. Then, we train the encoder and decoder of the model according to the two settings in GIM, as described in Section 3.4. We train our models on eight V100 GPUs with an initial learning rate (LR) of $6e^{-5}$ which is scheduled by the poly strategy with power 0.9 over 40 epochs. The optimizer is AdamW  [41] with epsilon $1e^{-8}$ weight decay $1e^{-2}$, and the batch size is 4 on each GPU.
Table 5: Benchmarking IMDL models for manipulation detection and localization for varying cross-generator generalization capability. The models are trained on the GIM-SD training set and tested on all the test sets to explore the generalization ability. The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported. $\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
Method GIM-SD GIM-SD(VOC) GIM-GLIDE GIM-DDNM
Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC
ManTranet[62] 73.12 43.18 80.18 63.18 27.37 72.74 58.21 24.53 74.61 39.52 16.8 58.31
MVSS-Net[12] 56.12 25.13 82.15 56.30 21.2 73.17 53.12 23.17 73.11 48.12 10.11 50.15
SPAN[25] 52.15 47.32 56.93 52.17 29.86 55.17 50.03 30.14 56.08 44.17 14.30 56.67
PSCC-Net[40] 58.24 48.36 87.29 54.79 39.36 83.80 51.11 29.60 79.76 40.52 4.40 48.52
[…]
SegFormer[63] 71.93 53.08 87.11 60.03 28.17 80.97 54.11 21.03 75.12 50.21 4.70 51.33
GIMFormer (Ours) 78.96 61.75 90.61 67.19 50.01 84.01 67.01 39.27 81.02 50.23 19.37 60.12
4.6 Comparison with state-of-the-art methods
We compare our methods with various state-of-the-art IMDL methods  [62, 12, 25, 40, 19, 58] on generative manipulation detection and location. In addition, the vanilla SegFormer (MiT-B2) [63] is also compared, since our method is based on that architecture. As shown in Table 4 and Table 5, these methods are tested on various settings to verify the performance and generalization. Note that some methods are not explicitly designed for image-level detection, in which case we use the maximum of the prediction map as the detection statistic. All methods are immersed in the same implementation details.
Cross-generator generalization capability comparison. Table 5 reports the generalization performance of all the methods mentioned. The results show that GIMFormer outperforms all other methods by a significant margin in both in-domain detection ability and cross-domain generalization ability. For in-domain detection ability, our method efficiently catches the subtle artifacts inherent in generative and accurately localizes them. Meanwhile, other methods may encounter confusion as they attempt to learn specific content, potentially leading to challenges in accurately detecting and localizing generative tampering patterns. For cross-domain generalization ability, GIMFormer achieves state-of-the-art performance. It demonstrates cross-detection capability in detecting manipulation using different generative models, as shown in the results of the GIM-GLIDE and GIM-DDNM testsets. Besides GIMFormer works well on Data generated by the same generator from different distributions, while existing methods are confused and have an obvious performance drop, as shown in the results of the GIM-SD(VOC) testset. The qualitative results for visual comparisons are illustrated in Figure 6.
Mix-generator comprehensive capability comparison Table 5 reports the comprehensive performance of all the methods mentioned. GIMFormer also outperforms all other methods on all the test sets, demonstrating its superior ability to identify generative tampering and its comprehensive performance. Existing methods have a low pixel-level F1 score on this benchmark, which means that tampering areas cannot be accurately identified. The qualitative results for visual comparisons are illustrated in Figure 5.
Figure 5: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods comprehensive capability. From top to bottom, we show examples from the authentic image, GIM-SD, GIM-GLIDE, GIM-DDNM
Figure 6: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods generalization capability. From top to bottom, we show examples from GIM-SD, GIM-GLIDE, GIM-DDNM
4.7 Ablation Analysis
Effectiveness of proposed module. We consider a simple baseline proposed in  [63] and gradually integrate new key components. Experiments are carried out on the GIM-GLIDE testset. The quantitative results are listed in Table 7. The result shows that ShadowTracer brings significant improvements to the vanilla baseline. With the MWAM, there is an increase of 6.29% in F1 and 2% in AUC, which indicates that the differential information at multiple scales is crucial for accurate tampering localization. The use of FSB to dynamically harvest complementary frequency and spatial cues improves performance, particularly in detection. The results verify that ShadowTracer, FSB and MWAM effectively improve the performance of the baseline model.
Table 6: Ablation results on GIM-GLIDE test dataset using different variants of GIMFormer. All detection Cls.Acc(%), localization AUC(%), and F1(%) scores are reported.
Variants Cls.Acc F1 AUC
Baseline 78.11 56.77 88.73
+ST 79.96 67.12 91.99
+ST+WMAM 80.19 73.41 94.01
+ST+WMAM+FSB 83.89 77.31 95.42
Table 7: Generalization Experiments of ShadowTracer in percentage (%). ST denotes ShadowTracer and ST(SD) denotes ShadowTracer trained on data generated by Stable-Diffusion.
Variants GIM-GLIDE GIM-DDNM
Cls.Acc F1 AUC Cls.Acc F1 AUC
GIMFormer w ST(SD) 83.1 78.1 95.4 77.4 58.8 89.4
GIMFormer w/o ST 80.0 68.9 91.3 73.1 51.3 86.1
Generalization of ShadowTracer. We initiate the training process of ShadowTracer using data generated by Stable Diffusion and subsequently hold its weights fixed. Following this pretraining phase, we proceed to train the backbone on both the GIM-GLIDE and GIM-DDNM trainsets, with and without the incorporation of ShadowTracer. The detection result Cls.Acc and location results in F1 and AUC are presented in Table 7, revealing that leveraging the pretrained ShadowTracer significantly enhances performance in cross-generator IMDL tasks. Notably, in the GIM-GLIDE dataset with the same manipulation type, leveraging ShadowTracer leads to a remarkable 9% enhancement in F1 and a 4% increase in AUC for location results. Additionally, the benefits extend to the detection aspect, manifesting as a 3% increase in accuracy. In the GIM-DDNM dataset with various manipulation types, there is a 7% enhancement in F1, a 3% increase in AUC for location results and a 4 % increase in accuracy.
5 Conclusion
We address the challenge of detecting and locating generative-based manipulation and provide a reliable database (GIM) for AIGC security. The proposed dataset leverages multiple mainstream generators and tampering methods to provide a variety of generative manipulation data. Additionally, We introduce GIMFormer, a novel transformer-based IMDL framework. ShadowTracer is designed to catch subtle artifacts in generative tampering. While the Frequency-Spatial Block gathers frequency and spatial information, the Multi Windowed Anomalous Modelling module captures pixel-level inconsistencies at multiple scales for fine-grained features. Extensive experiments demonstrate the superior performance of our model, which achieves SoTA results.
Limations. The GIM dataset shares classes with ImageNet and VOC, but it may not encompass future emerging objects due to the evolving variety in the real world. Fine-tuning pre-trained models on new object data can address this issue. Current research primarily centers on image manipulation. The emergence of realistic video generation models like SORA [5] presents fresh challenges as AI manipulation extends into videos. Future plans involve expanding research to address the complexities of video manipulation.
Supplementary Document
In this supplementary document, we include many details of our work: (1) The details of the proposed GIM Datset and compare with related work (Sec. 6) (2) The details of Architecture and implementation (Sec. 7). (3) The additional ablation analysis of GIMFormer(Sec. 8) (4) More Visualizations on GIM dataset (Sec. 9). (5)The societal impact of our work (Sec. 10). (6) The ethics statement of our work (Sec.11).
6 GIM Dataset
6.1 GIM Dataset Configuration
The specific quantities and details of the dataset are presented in Table 9, GIM has a total number of 1,140k generative manipulated images with their corresponding origin images. Moreover, the dataset can be split into subsets. GIM-SD-all utilizes the training set from Imagenet-1k [10] and employs the Stable Diffusion [53] generator for image manipulation dedicated to manipulation research. Through the analysis of the data scale, GIM-SD, GIM-GLIDE and GIM-DDNM individually select a random set of 100 classes from the Imagenet-1k training dataset and the entire test data to generate their respective training and test sets to benchmark the image manipulation detection and location (IMDL)methods. GIM-SD (VOC) is a cross-data-distribution test set constructed using the test set from the PASCAL VOC [14] dataset and Stable Diffusion techniques.
Table 8: Parameters for GIM degradation. Three degradation methods are employed with variable parameters to emulate real-world scenarios.
Post-processing Parameter
JPEG compression 75, 80, 90
Gaussian blur 3, 5
Downsampling 0.5, 0.67
Table 9: Basic configuration about the GIM Dataset. GIM has five subsets, leveraging various generators and data sources.
GIM
Tampreing
Type
Number of
Trainset
Number of
Testset
GIM-SD-all reprint 1890K -
GIM-SD reprint 180K 70K
GIM-GLIDE reprint 140K 70K
GIM-DDNM removal 125K 50K
GIM-SD(VOC) reprint - 4.5K
To simulate real-world scenarios and comprehensively evaluate the IMDL methods, GIM incorporates three degradation methods (JPEG compression, Gaussian blur, and downsampling). Table 9 presents the parameters for each degradation group, which are randomly chosen.
The GIM data format consists of JPG images paired with PNG labels. Filenames ending with "_f" indicate tampered images. The masks segmented by SAM [33] serve as ground truth for manipulation detection, containing two labels denoted by 0 and 255, representing the original and manipulated categories, respectively.
6.2 Dataset Comparison.
GIM focuses on generative local manipulation in natural images and provides manipulated region labels, specifically for image manipulation detection and localization tasks. This distinguishes GIM from the existing datasets focused on faces or limited to detection (classification). As shown in Table 10, we compare several local image manipulation datasets. The existing image forensic datasets are primarily constructed using traditional manipulation techniques. These datasets commonly employ manual or random manipulation to generate data. AutoSplice [29], HiFi-IFDL[20], CoCoGlide  [19] dataset utilizes generative algorithms for image manipulation. However, they rely on one type of manipulation and suffer from a small data volume.
In contrast, GIM achieves image manipulation by employing generative methods for localized image alteration. It encompasses a variety of generators and manipulation methods. Simultaneously, it stands as a large-scale comprehensive dataset for the IMDL community.
Table 10: Summary of previous image manipulation datasets and GIM. We showcase the number of data entries within each dataset and the manipulation techniques they encompass. *denotes that HiFi-IFDL includes entirely synthesized images
Dataset Num. Images Handcraft Manipulations AIGC Manipulations
# Authetic Image # Forged Image Splicing Copy-move Removal Removal Generation
Columbia Gray [43] 933 912
Columbia Color [24] 183 180
MICC-F2000 [1] 1,300 700
VIPP Synth.  [1] 4,800 4,800
CASIA V1.0  [13] 800 921
CASIA V2.0  [13] 7,200 5,123
Wild Web [66] 90 9,657
NC2016  [17] 560 564
NC2017  [17] 2,667 1,410
MFC2018  [17] 14,156 3,265
MFC2019  [17] 10,279 5,750
PS-Battles [22] 11,142 102,028
DEFACTO  [42] 0 229,000
IMD2020 [45] 35,000 35,000
SP COCO  [36] 0 200,000
CM COCO  [36] 0 200,000
CM RAISE  [36] 0 200,000
HiFi-IFDL[20] - 1,000,00*
CoCoGlide [19] 0 512
AutoSplice  [29] 3,621 2,273
GIM 1,140,000 1,140,000
7 Implementation Details
7.1 Architecture
[…]
7.2 Training Process
We conduct our experiments with PyTorch 1.7.0. All models are trained on a node with 8 or 4 V100 GPUs.
For the training of the ShadowTracer network $g_{\phi}$, We randomly selected 20,000 authentic images from ImageNet and correspondingly generated manipulated images. During comprehensive performance verification, all three generators (Stable Diffusion [52], GLIDE [44], DDNM [59])are used, whereas for generalization verification, only Stable Diffusion is employed. Throughout the training, the alpha blending parameter ranges between 0.5 and 1.0, randomly producing blended images subjected to three degradation types (JPEG compression, downsampling, Gaussian blur). The network is trained on 40$\times$40 pixel patches randomly sampled from the dataset. Training is conducted for roughly 300,000 iterations with a batch size of 64. An Adam [31] optimizer is employed, initialized with a learning rate of 0.001.
For the training of the GIMFormer backbone network. The input image is cropped to 512 $\times$ 512 during training. We train our models for 40 epochs on the GIM dataset. The batch size is 4 on each of the GPUs. The images are augmented by random resize with a ratio of 0.5–2.0, random horizontal flipping.
8 Additional Ablation Analysis
The robustness of GIMFormer. To evaluate the robustness of GIMFormer, we sample 5000 images from the original three generator test sets without degradations. The model is trained on the mixed-generator dataset. The model is trained on the mixed-generator dataset. The experimental results are presented in Table 11, GIMFormer demonstrates robustness against various distortion techniques. This resilience underscores its ability to maintain stable performance in the face of various challenges posed by image distortions.
Table 11: Robustness experiment of pixel-level Manipulation Localization F1(%) with various distortions
No Dis-
tortion
Cmp
(q=90)
Cmp
(q=80)
Cmp
(q=75)
Blur
(k=3)
Blur
(k=5)
Downsample
Downsample
GIMFormer 61.7 60.3 60.1 59.9 58.7 58.7 61.0 60.8
The effect of the number of windows (N). We conduct a set of ablation experiments to study the performance of the MWAM Module. To ensure fair comparisons, all experiments differ from each other only in the Windows setting. Experiments are carried out on the GIM-SD trainset and testset. As shown in Table 12, there is an overall incremental trend in the tampering location performance as the number of windows increases, while the impact of size variations remains relatively minor. For the sake of efficiency, we stop analyzing more windows. The results indicate that a favorable balance between accuracy and efficiency is achieved when $N$=3.
Table 12: Ablation results on GIM-SD test dataset using different variants of GIMFormer, All detection Cls.Acc(%) and localization AUC(%) and F1(%) scores are reported. Where $D_{a/m}^{k}$ represents the number and dimensions of windows utilized in either the average or maximum branches. We investigate the window counts $N$ of 0, 1, 2,3 as well as the impact of only one branch on the module. In this table, the window sizes of the first layer are used to annotate, with subsequent layers decreasing in size.
MWAM Variants Cls.Acc F1 AUC
[…]
{$D_{a\&k}^{11}$} 74.33 58.34 89.65
{$D_{a\&k}^{11},D_{a\&k}^{21}$} 76.61 60.03 89.92
{$D_{a\&k}^{11},D_{a\&k}^{21},D_{a\&k}^{31}$} 78.96 61.75 90.61
{$D_{a\&k}^{17},D_{a\&k}^{29},D_{a\&k}^{41}$} 77.33 61.17 90.11
{$D_{a}^{11},D_{a}^{21},D_{a}^{31}$} 77.95 60.77 90.03
{$D_{m}^{11},D_{m}^{21},D_{m}^{31}$} 77.83 59.13 89.53
{$D_{a\&k}^{11},D_{a\&k}^{21},D_{a\&k}^{31},D_{a\&k}^{41}$} 78.11 61.15 90.71
9 More Visualizations on GIM
In this section, we present additional visualizations of the GIM dataset in Figure  7,  8, and  9, showcasing the data visualization from GIM-SD, GIM-GLIDE, and GIM-DDNM respectively.
10 Societal Impact
Our research yields a positive societal impact on the community by focusing on addressing the challenge of detecting and locating generative-based manipulations. We introduce a reliable database, GIM, aimed at enhancing the security of AI-generated content (AIGC). This database facilitates the training of Image Manipulation Detection and Localization (IMDL)methods on GIM, allowing them to generalize across various scenarios. Consequently, our algorithm GIMFormer fosters increased trust among the general public in our society regarding media content. GIM and GIMFormer are beneficial for digital media forensics, especially generative manipulation with real-world degradations.
11 Ethics Statement
Our GIM is based on ImageNet and VOC. No additional personally identifiable information or sensitive personally identifiable information is introduced during the production of fake images in the GIM dataset. During the dataset production, we do not introduce extra information containing exacerbated bias against people of a certain gender, race, sexuality, or who have other protected characteristics. The ethical issues in the ImageNet and VOC datasets have been discussed in previous works. Crawford et al.  [8] discuss issues with ImageNet. The first issue is the political nature of all taxonomies or classification systems, where terms like "male" and "female" are considered "natural," while "hermaphrodite" is offensively placed within the branch of Person > Sensualist > Bisexual alongside "pseudohermaphrodite" and "switch hitter" categories. The second issue concerns offensive images of real people, while the third is the use of people’s photos without their consent by ImageNet creators.
Figure 7: Additional visualizations of GIM-SD. From left to right, the sequence is as follows: original image, tampered area, generative manipulated image, original image, tampered area, generative manipulated image.
Figure 8: Additional visualizations of GIM-GLIDE. From left to right, the sequence is as follows: original image, tampered area, generative manipulated image, original image, tampered area, generative manipulated image.
Figure 9: Additional visualizations of GIM-DDNM. From left to right, the sequence is as follows: original image, tampered area, generative manipulated image, original image, tampered area, generative manipulated image.

### v2

@@
 The extraordinary ability of generative models emerges as a new trend in image editing and generating realistic images, posing a serious threat to the trustworthiness of multimedia data and driving the research of image manipulation detection and location (IMDL).
-However, the lack of a large-scale data foundation makes IMDL task unattainable.
-In this paper, a local manipulation pipeline is designed, incorporating the powerful SAM, ChatGPT and generative models.
-Upon this basis, We propose the GIM dataset, which has the following advantages: 1) Large scale, including over one million pairs of AI-manipulated images and real images.
-2) Rich Image Content, encompassing a broad range of image classes 3) Diverse Generative Manipulation, manipulated images with state-of-the-art generators and various manipulation tasks.
+However, the lack of a large-scale data foundation makes the IMDL task unattainable.
+In this paper, we build a local manipulation data generation pipeline that integrates the powerful capabilities of SAM, LLM, and generative models.
+Upon this basis, we propose the GIM dataset, which has the following advantages: 1) Large scale, GIM includes over one million pairs of AI-manipulated images and real images.
+2) Rich image content, GIM encompasses a broad range of image classes.
+3) Diverse generative manipulation, the images are manipulated images with state-of-the-art generators and various manipulation tasks.
 The aforementioned advantages allow for a more comprehensive evaluation of IMDL methods, extending their applicability to diverse images.
-We introduce two benchmark settings to evaluate the generalization capability and comprehensive performance of baseline methods.
-In addition, we propose a novel IMDL framework, termed GIMFormer, which consists of a ShadowTracer, Frequency-Spatial Block (FSB), and a Multi-window Anomalous Modelling (MWAM) Module.
-Extensive experiments on the GIM demonstrate that GIMFormer surpasses previous state-of-the-art works significantly on two different benchmarks.
-Code page: GIM.
-Keywords:
-Dataset and Benchmark Image Manipulation Detection and Location Forgery Traces Extraction Multi-scale Feature Learning
-11footnotetext: * These authors contributed equally to this research, which was done during Yirui Chen and Quan Zhang’s internship at Huawei Noah’s Ark Lab. ‡ Project lead † Corresponding authors
-1 Introduction
-Figure 1: Example image from the GIM dataset.
+We introduce the GIM benchmark with two settings to evaluate existing IMDL methods.
+In addition, we propose a novel IMDL framework, termed GIMFormer, which consists of a ShadowTracer, Frequency-Spatial block (FSB), and a Multi-Window Anomalous Modeling (MWAM) module.
+Extensive experiments on the GIM demonstrate that GIMFormer surpasses the previous state-of-the-art approach on two different benchmarks.
+Introduction
+Figure 1: Example images from the GIM dataset.
 Our dataset includes images manipulated by three state-of-the-art generators: Stable-Diffusion, GLIDE, and DDNM.
-The first column shows authentic images, while the second column presents forgery masks.
-The third column displays forged images.
-Images are one of the most essential media for information transmission in modern society, which is widely spread on public platforms such as news and social media.
-With the rapid advancement of generative methods [11, 52, 16, 4], such natural information can be more easily tampered with for specific purposes such as tampering with an object or person.
-The image of this class is particularly convincing since its visual comprehensibility, leads to serious information security risks in social areas.
-For instance, an AI-generated image of black smoke billowing from what appeared to be a government building near the Pentagon set off investor fears, sending stocks tumbling [46].
-Therefore, it is of utmost urgency to develop effective methods to examine whether an image is modified by generative models, meanwhile identifying the exact location of the tampering.
-However, the traditional image manipulation detection and location(IMDL) dataset [55, 13, 60] overlook the powerful generative models and generative IMDL dataset are limited with scale, thus not sufficient to comprehensively evaluate the performance of IMDL methods and benefits the IMDL community.
-To this end, we propose a million-level generative-based IMDL dataset, termed GIM, to provide a reliable database for AI Generated Content (AIGC) security.
-GIM leverage the diffusion model [44, 52, 59] and SAM [32, 51], with ImageNet [10] and VOC [14] as the input.
-SAM is leveraged to locate the tampering region, and diffusion models paint the reasonable content in the tampering area.
-The excellent diffusion model ensures dataset fidelity and rich categories guarantee the diversity of images in GIM.
-GIM contains a total of over one million generative tampered images.
-Based on this, the IMDL methods are evaluated and benchmarked.
+Three columns display authentic images, manipulation masks and manipulated images.
+Images are one of the most essential media for information transmission in modern society and they are widely spread on public platforms such as news and social media.
+With the rapid advancement of generative methods (Dhariwal and Nichol 2021; Rombach et al.
+2022), such natural information can be more easily manipulated for specific purposes such as tampering with an object or person.
+The image of this class is particularly convincing since its visual comprehensibility, leading to serious information security risks in social areas.
+Therefore, it is of utmost urgency to develop methods to detect whether an image is modified by generative models and identify the exact location of the manipulation.
+However, traditional image manipulation detection and location (IMDL) datasets (Dong, Wang, and Tan 2013; Wen et al.
+2016) overlook the powerful generative models and generative IMDL (Jia et al.
+2023; Guillaro et al.
+2023) datasets are limited with scale, it is thus not sufficient to comprehensively evaluate the performance of IMDL methods and benefit the IMDL community.
+To this end, we propose a million-level generative-based IMDL dataset, termed GIM dataset, to provide a reliable database for AI Generated Content (AIGC) security.
+GIM leverages the generative models (Ho, Jain, and Abbeel 2020; Yin et al.
+2025) and SAM (Kirillov et al.
+2023a), with the images in ImageNet (Deng et al.
+2009) and VOC (Everingham et al.
+2010) as the input.
+SAM is utilized to locate the tampering region, and generative models paint the reasonable content in the tampering area.
+GIM contains a total of over one million generative manipulated images.
 To develop an appropriate benchmark scale, we explore the impact of different amounts of training manipulated data.
-The final benchmark contains about 300k manipulated images with their tampering masks.
-To simulate the real situation, we investigate the effect of degradation and apply the images to the random degradation method.
-Two benchmark settings are designed to evaluate the comprehensive and generalization.
-In addition, we recreated existing IMDL methods in a fair manner, providing the basis for future development.
-Overall, as a generative IMDL benchmark, GIM possesses the following advantages: 1) GIM has a large and reasonable data scale, including rich image categories and contents.
+The final benchmark contains about 320k manipulated images with their tampering masks for training and testing.
+To simulate the real situation, we investigate the effect of degradation and apply random degradation to these benchmark images.
+Based on this GIM benchmark, the IMDL methods are evaluated and benchmarked.
+Overall, GIM possesses the following advantages: 1) GIM has a large and reasonable data scale, including rich image categories and contents.
 2) GIM contains various generative manipulation models and tasks.
-3) GIM proposes two settings for verifying generalization capability and comprehensive performance in existing methods.
-The existing methods emphasize traditional image manipulations also known as “cheapfakes”.
-Generative manipulations introduce lethal alterations in content with no apparent frequency or structural inconsistency.
+3) GIM proposes two settings for verifying the performance and generalization ability of IMDL methods.
+Existing methods emphasize traditional image manipulations also known as “cheapfakes”.
+However, generative manipulations introduce lethal alterations in content with no apparent frequency or structural inconsistency.
 To address the above issue, we introduce GIMFormer, a transformer-based framework for generative IMDL.
-The ShadowTracer is designed to embed the nuanced artifacts inherent in generative tampering and serves as prior information.
-The Frequency-Spatial Block harvests the manipulation clues in frequency and spatial domains.
+The ShadowTracer is designed to embed the nuanced artifacts inherent in generative tampering and serves as the prior information.
+The Frequency-Spatial Block captures the manipulation clues in the frequency and spatial domains.
 Furthermore, the Multi Windowed Anomalous Modelling module captures local inconsistencies at different scales to refine the features.
-By doing so, our model extracts features from both the RGB and learned tampering traces map, captures details from the frequency domain and spatial domain, and models inconsistencies at different scales for precise manipulation detection and location.
+GIMFormer extracts features from both the RGB and learned tampering trace maps, capturing details from both the frequency and spatial domains while modeling inconsistencies at different scales for precise manipulation detection and localization.
 We conduct experiments on our proposed GIM.
-Both the qualitative and quantitative results demonstrate that GIMFormer can significantly outperform previous state-of-the-art methods in terms of both comprehensive capability and generalization ability.
+Both the qualitative and quantitative results demonstrate that GIMFormer can outperform previous state-of-the-art methods.
 In summary, our main contributions are as follows:
-  • We build a local manipulation pipeline and construct a comprehensive large-scale dataset with various SOTA generative models and manipulation tasks to further facilitate the research on the IMDL task.
-  • We investigate the impact of data scales and degradation on IMDL tasks to select an appropriate scale and simulate real-world scenarios.
-Based on this, we construct a reasonable benchmark and two settings to evaluate the generalization and comprehensive capabilities of IMDL methods.
-  • We propose GIMFormer, a generative IMDL framework consisting of ShadowTracer to embed subtle traces, Frequency-Spatial Block to extract forgery features in the frequency and spatial domains and Multi-Windowed Anomalous Modelling Module to capture pixel-level inconsistencies at multi scales.
-  • Extensive experiments show that GIMFormer achieves state-of-the-art IMDL performance in terms of generalization and comprehensive capabilities.
-2 Related Work
-2.1 Image Forensic Datasets
-The research community has dedicated significant efforts to establishing robust datasets for image forensics.
-Early datasets primarily focus on one type of manipulation.
-The Columbia dataset  [43, 24] is the pioneering dataset focusing on splicing forgeries.
-CASIA v1.0 and CASIA v2.0  [13] are the first to incorporate multiple types of manipulations within a single dataset, with forged images manually crafted using Adobe Photoshop.
-The Wild Web [66] dataset collects forged images from the Internet, significantly surpassing the scale of previous datasets.
-NIST [17] Datasets provide an extensive collection of datasets serving as a crucial standard for assessing media tampering detection methods.
-DEFACTO [42]is automatically generated from the Microsoft COCO dataset, encompassing four categories of forgeries: Copy and Move, Splicing, Object Removal, and Morphing.
-IMD2020 [45] provides a range of locally manipulated images generated through manual operations or random slicing, while also collecting online images without apparent manipulation traces.
-Recently, HiFi-IFDL [20] construct a hierarchical fine-grained dataset containing some representative forgery methods.
-AutoSplice [29] leverage the capabilities of large-scale language-image models like DALL-E2  [48] to facilitate automatic image editing and generation.
-CocoGlide [19] contains 512 images using the GLIDE diffusion model.
-Some datasets focused exclusively on facial manipulations [54, 18, 9] or entirely synthesized images[71, 56, 3].
-However, these datasets have certain limitations, such as small data sizes and limited manipulation techniques or tasks.
-Recent advances in generative models [44, 23, 52] have demonstrated remarkable abilities in generative-based manipulation.
-These developments have led to the emergence of open-source projects [51, 65, 15] for manipulation.
-Leveraging the power of these models, we introduce GIM, a large-scale generative-based manipulation dataset.
+  • We build a data generation pipeline and construct a large-scale dataset for IMDL tasks.
+  • We investigate the impact of data scales and degradation, constructing a comprehensive benchmark with two evaluation settings for IMDL methods evaluation.
+  • We propose a framework named GIMFormer for generative IMDL task. Extensive experiments demonstrate that GIMFormer achieves state-of-the-art performance.
+Related Work
+Image Forensic Datasets
+Early datasets (Ng, Chang, and Sun 2004) primarily focus on one type of manipulation.
+CASIA  (Dong, Wang, and Tan 2013) first incorporates multiple types of manipulations with forged images manually crafted using Adobe Photoshop.
+The Wild Web dataset (Zampoglou, Papadopoulos, and Kompatsiaris 2015) collects forged images from the Internet, surpassing previous datasets in the scale.
+NIST (Guan et al.
+2019) dataset provides an extensive collection of datasets serving as a crucial standard for assessing media tampering detection methods.
+IMD2020 (Novozamsky, Mahdian, and Saic 2020) provides locally manipulated images generated through manual operations or random slicing and online images without obvious manipulation traces.
+Recently, HiFi-IFDL (Guo et al.
+2023) constructs a hierarchical fine-grained dataset containing some representative forgery methods.
+AutoSplice (Jia et al.
+2023) leverages large-scale language-image models like DALL-E2  (Ramesh et al.
+2021) to facilitate automatic image editing and generation.
+CocoGlide (Guillaro et al.
+2023) contains images manipulated by the diffusion model.
+Besides, other datasets focus on facial manipulations (Rossler et al.
+2019; Guarnera et al.
+2022) or entirely synthesized images(Zhu et al.
+2023; Yan et al.
+2024; Verdoliva and Cozzolino 2022).
+The datasets mentioned above have limitations such as small data sizes and limited manipulation techniques.
+Recent advances in generative models have demonstrated remarkable manipulation abilities.
+Leveraging these models, we introduce GIM, a large-scale dataset that incorporates various recent generative manipulation techniques.
 Figure 2: An overview of the dataset generation.
-With the input of the original image and the desired region of interest, classification attribution or mouse input as a User query, the manipulation mask is extracted by the zero-shot segmentation model SAM.
-Besides, interacting with ChatGPT, we organize the tampering prompt by combining replacement classes.
-The final generations are produced by generative models with the input of the original image, tampering mask and tampering prompts.
-2.2 Image Manipulation Detection and Localization
-Early studies on natural image manipulation localization mainly focus on detecting a specific type of manipulation [7, 27, 34, 50].
-Due to the exact manipulation type is unknown in real-world scenarios, most state-of-the-art methods [40, 19, 58, 64, 38, 28, 61] primarily concentrate on general manipulation.
-RGB-N [70] propose a dual-stream network for manipulation detection.
-One stream focuses on extracting RGB features and the other leverages noise features to model the disparities.
-MVSS-Net [12] utilizes a two-stream CNN to extract noise features and employs the Dual Attention Module to merge the output of the two-stream CNN.
-PSCC-Net  [40] extracts hierarchical features with a top-down path and detects whether the input image has been manipulated using a bottom-up path.
-ObjectFormer [58] uses object prototypes to model object-level consistencies and find patch-level inconsistencies to detect the manipulation.
-However, these methods focus on "cheapfake" detection and encounter challenges when applied to generative manipulation due to the tampering traces are subtle and the lack of an understanding of generative manipulation patterns.
-Certain works [26, 6, 9, 39, 2] focus on localizing manipulations using generative models, yet these methods primarily concentrate on human faces.
-In this work, we leverage a deep network to embed the subtle artifacts inherent in generative tampering as a prior trace map.
-we then employ a dual network to fuse the trace map and RGB image, combining frequency and spatial information and capturing pixel-level inconsistencies at multiple scales to detect and locate the manipulation.
-Table 1: Summary of previous image manipulation datasets and GIM.
-We showcase the number of data entries within each dataset, the image content, image sizes and the manipulation techniques they encompass.
-$\star$ denotes that HiFi-IFDL is composed of multiple datasets, including entirely synthesized images, traditionally manipulated images and fake face images.
+Given the original image and a user query (classification attribution or mouse input), the manipulation mask is extracted using SAM.
+Tampering prompts are then organized with LLM by combining replacement classes.
+The final generations are produced by generative models with the image, tampering mask and prompts.
+Image Manipulation Detection and Localization
+Early studies on natural image manipulation localization mainly focus on detecting a specific type of manipulation (Cozzolino, Poggi, and Verdoliva 2015; Yin et al.
+2023).
+Due to the exact manipulation type is unknown in real-world scenarios, most state-of-the-art methods (Liu et al.
+2022; Ying et al.
+2023) primarily concentrate on general manipulation.
+MVSS-Net (Dong et al.
+2022) uses a two-stream CNN to extract noise features and employs dual attention to fuse their outputs.
+PSCC-Net (Liu et al.
+2022) extracts hierarchical features in a top-down manner and it detects manipulations in a bottom-up manner.
+Motivated by the power of transformer(Zhang et al.
+2024; Zhang and Qi 2024; Lu et al.
+2024a, b; Chen et al.
+2023), ObjectFormer (Wang et al.
+2022) uses object prototypes to model object-level consistencies and find patch-level inconsistencies to detect the manipulation.
+However, these methods focus on “cheapfake” detection and encounter challenges when applied to generative manipulation.
+Certain works (Huang et al.
+2022; Chai et al.
+2020; Liu et al.
+2024) localize manipulations using generative models but mainly focus on human faces.
+In this work, we focus on the generative IMDL task.
+We leverage a deep network to embed the subtle artifacts inherent in generative tampering as a prior trace map.
+A dual network fuses the trace map and RGB image, combining frequency and spatial information and capturing pixel-level inconsistencies to detect and locate the manipulation.
+Dataset and Benchmark Construction
+In this section, We propose an automatic pipeline to generate manipulated images from unannotated data.
+Leveraging this pipeline, we construct the comprehensive large-scale GIM dataset.
+To build a reasonable benchmark, we conduct preliminary experiments focusing on two key aspects: data scale and image degradation.
+First, we investigate the impact of the training data scale to determine an appropriate size for the GIM benchmark.
+Second, to reveal real-world scenarios, the manipulated and original images are subjected to three random degradations.
+Finally, the GIM benchmark comprises over 320k manipulated images paired with authentic images for algorithm training and evaluation.
+The example images are shown in Figure 1 with their original images and tampering masks.
+Finally, we introduce the criteria and settings used to evaluate the IMDL methods.
 Dataset Image Content Image Size Num. Images Manipulations Category
-# Authetic Image # Forged Image Traditional Gen.Reprint Gen.Removal
-FaceForensics++ [54] Face 480p-1080p 1000 4000
-DFDC [9] Face 240p-2160p 19,154 100,000
-DeeperForensics [30] Face 1080p 50,000 10,000
-Columbia Gray [43] General 128 $\times$ 128 933 912
-CASIA V2.0  [13] General 320 $\times$ 240-800 $\times$ 600 7,200 5,123
-IMD2020 [45] General 1062 $\times$ 866 35,000 35,000
-Coverage [60] General 400 $\times$ 486 100 100
-AutoSplice  [29] General 256 $\times$ 256 - 4232 $\times$ 4232 3,621 2,273
-HiFi-IFDL[20] General - - 1,000,000$\star$
-CocoGlide [19] General 256$\times$256 - 512
+# Authentic Image # Forged Image Traditional Gen.Reprint Gen.Removal
+FaceForensics++ (Rossler et al. 2019) Face 480p-1080p 1,000 4,000
+DFDC (Dang et al. 2020) Face 240p-2160p 19,154 100,000
+DeeperForensics (Jiang et al. 2020) Face 1080p 50,000 10,000
+Columbia Gray (Ng, Chang, and Sun 2004) General 128 $\times$ 128 933 912
+CASIA V2.0  (Dong, Wang, and Tan 2013) General 320 $\times$ 240-800 $\times$ 600 7,200 5,123
+IMD2020 (Novozamsky, Mahdian, and Saic 2020) General 1062 $\times$ 866 35,000 35,000
+Coverage (Wen et al. 2016) General 400 $\times$ 486 100 100
+AutoSplice  (Jia et al. 2023) General 256 $\times$ 256 - 4232 $\times$ 4232 3,621 2,273
+HiFi-IFDL (Guo et al. 2023) General - - 1,000,000 (200,000)
+CocoGlide (Guillaro et al. 2023) General 256$\times$256 - 512
 GIM General 64$\times$64-6000$\times$3904 1,140,000 1,140,000
-3 Dataset and Benchmark Construction
-In Section 3.1, we propose an automatic data synthesis pipeline to produce generative manipulated images efficiently with unlabeled images.
-Leveraging this pipeline, we construct a comprehensive large-scale dataset.
-To build a reasonable scale benchmark for evaluating the IMDL methods, we delve into the impact of training data scale in Section 3.2.
-To emulate real-world conditions and establish an objective benchmark, the manipulated images and original authentic images are subjected to three random degradations (JPEG compression, downsampling, and Gaussian blur), as detailed in Section 3.3).
-GIM benchmark comprises over 320k manipulated images paired with authentic images (100 distinct labels in ImageNet for each generative model) for algorithm evaluation.
-The generated images are shown in Figure 1 with their original images and tampering masks.
-The specific composition of GIM can be found in the Supplementary Materials.
-In Section 3.4, we introduce the criteria and settings used to evaluate the performance of IMDL methods.
-3.1 Generation and Details of the dataset
-To construct a generative-based IMDL dataset, a large-scale natural dataset is urgently desired.
-ImageNet [10] and VOC [14] are awesome datasets, chosen as the starting point for our research due to the following advantages: 1) Large and diverse.
-2) Wide category coverage.
-To sum up, we regard these datasets as the database for generative manipulation.
-In addition to possessing various scenarios, the proposed dataset utilizes a wealth of tampering methods.
-Specifically, we reprint the target class by generative inpainting [52, 44] method or remove the destination region using the model [59].Benefiting from open-source projects [33, 51], we develop our data generation pipeline.
+Table 1: Summary of image manipulation datasets.
+Image content, image sizes, and manipulation techniques are reported.
+HIFI-IFDL includes mainly entirely synthesized images with a small portion of traditional manipulation or face image manipulation.
+Data Generation Pipeline
+Benefiting from open-source projects (Kirillov et al.
+2023b; Ren et al.
+2024), we develop our data generation pipeline.
 Figure 2 illustrates the overall process.
-Firstly, with the classification attribute or User query, the local manipulation mask is extracted by the zero-shot segmentation network [33].
-For reprint tampering, the image category is embedded in the replacement prompt and interacts with ChatGPT, which returns an approximate category.
+Our pipeline includes two types of manipulation: reprinting the target class using a generative inpainting method Stable Diffusion (Rombach et al.
+2022) and GLIDE  (Nichol et al.
+2021) or removing the destination region using a specific removal method DDNM (Wang, Yu, and Zhang 2022).
+Given an arbitrary image, the process begins by extracting the local manipulation mask using a zero-shot segmentation network SAM, guided by either classification attributes or user queries.
+For reprint tampering, the image category is embedded in the replacement prompt and interacts with LLM (Achiam et al.
+2023), which returns an approximate category.
 The approximate category is then embedded into the inpainting prompt.
-Combined with the original image, manipulation mask, and inpainting prompt, generative models generate the reprint generative tampered result.
-For removal tampering, only the original image and manipulation mask are required for the generated model.
-The GIM Dataset utilizes the entire ImageNet dataset and the Stable-Diffusion generator to construct a comprehensive dataset, providing a reliable database and laying the foundation for the research below.
-3.2 Analysis of Benchmark Scale
-With the data generation pipeline described above, we are qualified to generate millions, tens of millions, or even billions of data.
-Nonetheless, blindly increasing the amount of data does not improve the algorithm performance, but may lead to data redundancy.
-Therefore, taking data generated by the Stable Diffusion as an example, we explore the influence of data volume and category volume on baseline classification [21] and segmentation algorithms [63].
-Training sets are the different scales of the training subset, and the validation is the same test set.
+By combing with the original image, manipulation mask and inpainting prompt, generative models generate the reprint tampered result.
+For removal tampering, only the original image and manipulation mask are required for the model.
+The GIM dataset utilizes images from the ImageNet dataset  (Deng et al.
+2009) and VOC dataset  (Everingham et al.
+2010).
+These two datasets contain a large number of diverse images with wide category coverage, providing a reliable database and laying the foundation for future research.
+Analysis of the Benchmark Scale
+With the data generation pipeline, we are qualified to generate large amounts of manipulated images.
+Nonetheless, blindly increasing the amount of data does not improve the algorithm performance, while it may lead to data redundancy.
+Therefore, taking data generated by the Stable Diffusion as an example, we explore the influence of data volume and category volume on baseline classification (He et al.
+2016) and segmentation algorithms (Xie et al.
+2021).
+Training sets vary in the scale, while the validation uses the same test set.
 As shown in Table 2, the metrics of classification and segmentation are improved as the dataset scale increases.
-When the scale reaches 180K, the algorithm performance is almost saturated.
-No matter whether the image class or data scale is added, the metrics are stagnant.
+When the scale reaches 180K, the algorithm performance almost saturates.
+All the metrics remain almost unchanged with either the image class number or the data scale increased.
 Experiments demonstrate that increasing the amount of data or category brings negligible benefits when the data tends to be saturated.
-According to the analysis above, for the training data, the GIM benchmark selects 100 different labels from ImageNet for each generation model to create tampered images.
-For the test data, the GIM benchmark uses the entire test dataset for each generation model.
-Table 2: Results of classification and segmentation on different scale subsets. Classification uses Resnet-50 and segmentation uses SegFormer-b0.
-Total Num. of Image Image Classes Image per Class Metrics
-[…]
-2,800 10 280 0.6306 0.2550 0.7950
-28,000 100 280 0.7480 0.3247 0.7995
-180,000 100 1800 0.9129 0.5289 0.8699
-360,000 200 1800 0.9131 0.5291 0.8701
-500,000 500 1000 0.9131 0.5291 0.8703
-Table 3: Influence of different degradation method on classification (Resnet-50) and segmentation (SegFormer-b0) .
-Degradation Metrics
+According to the analysis, the GIM benchmark uses 100 labels from ImageNet to generate tampered images for training and employs all the test sets from ImageNet and VOC for evaluation.
+Total Num. of Image Image Classes Image per Class Metrics(%)
+Cls.Acc Seg.F1 Seg.AUC
+2,800 10 280 63.1 25.5 79.5
+28,000 100 280 74.8 32.5 79.9
+180,000 100 1800 91.3 52.9 86.9
+360,000 200 1800 91.3 52.9 87.0
+500,000 500 1000 91.3 52.9 87.0
+Table 2: Dataset Scale Experiment: Effect of Dataset Scale on the Performance of Base Models (SegFormer-b0, ResNet-50).
+Degradation Metrics(%)
 Cls.Acc Seg.F1
-- 91.29 52.89
-JPEG 83.10 45.74
-Gaussian Blur 90.94 40.14
-Downsample 88.34 37.12
-3.3 Post-processing of Degradation Method
-After being spread on the Internet, images will encounter various post-processings, such as JPEG compression, downsampling, Gaussian blur, etc.
+- 91.3 52.9
+JPEG compression 83.1 45.7
+Gaussian blur 90.9 40.1
+Downsample 88.3 37.1
+Table 3: Degradation Experiment: Impact of Degradation on Base Models (SegFormer-b0, ResNet-50).
+Post-processing of Degradation Method
+When uploaded to the Internet, images will encounter various post-processing.
 These transformations can pose challenges for image forensics methods.
-Figure 3.2 investigates the impact of degradation.
-The baseline models are trained on the clean data generated by the Stable-Diffusion training set and tested on the same test set with a specific degradation method.
-Experimental results denote that these degradation methods make identification more difficult.
-To shorten the distance from reality and build an objective benchmark, the random degradation method (JPEG compression, downsampling, Gaussian blur) is performed on the dataset.
-The details of the parameters can be found in Supplementary Materials.
-3.4 Benchmark Settings
-Dataset: The datasets used for the research of generative IMDL tasks are lacking.
-This makes direct comparisons and accurate measurement difficult.
-We use three cross-generator datasets GIM-SD (data from ImageNet manipulation by Stable Diffusion, similarly hereinafter), GIM-GLIDE, GIM-DDNM and a cross-distribution dataset GIM-VOC(data from VOC manipulation by Stable Diffusion) to present benchmarking results.
-Metrics: We evaluate the performance of the proposed method on both the image manipulation detection task and localization task.
-For classification results, we use Accuracy (Cls.acc) as our evaluation metric, while for localization, the pixel-level AUC and F1 score on manipulation masks are adopted.
-Since binary masks and detection scores are required to compute F1 scores, we adopt the Equal Error Rate (EER) threshold to binarize them.
-Settings: Two settings are proposed to verify the comprehensive and generalization performance of the algorithms.
-In the cross-generator generalization setting, the models are trained on the GIM-SD training set and tested on the GIM-GLIDE, GIM-DDNM and GIM-VOC test set to explore the generalization IMDL performance, as shown in Table 5.
-In the mix-generator comprehensive setting, the models are jointly trained on the GIM-SD, GIM-GLIDE and GIM-DDNM training set and tested on the correspondence test dataset respectively to evaluate the comprehensive IMDL performance, as shown in Table 4.
-4 Method
-[…]
+Table 3 investigates the impact of degradation.
+The baseline models are trained on the clean data and tested on the test set with a single degradation.
+Experimental results show that these degradations make the identification difficult.
+To reveal the real-world scenarios, random degradation (JPEG compression, downsampling and Gaussian blur) is performed on the dataset.
+Benchmark Settings
+Benchmark Description: GIM consists of four subsets.
+GIM-SD, GIM-GLIDE and GIM-DDNM which contain data from ImageNet manipulated by Stable Diffusion, GLIDE and DDNM, respectively.
+Additionally, there is a cross-distribution subset GIM-VOC which contains data from VOC manipulated by Stable Diffusion.
+Metrics: We evaluate the performance of the IMDL method on both the image manipulation detection task and localization task.
+For the detection task, we use accuracy (Cls.acc) as our evaluation metric.
+For the localization task, the pixel-level AUC and F1 score are adopted.
+Settings: Two settings are proposed to evaluate the performance and generalization.
+In the mix-generator setting, the models are jointly trained on the GIM-SD, GIM-GLIDE and GIM-DDNM training set and tested on the correspondence test dataset respectively to evaluate the performance.
+In the cross-generator setting, the models are trained on the GIM-SD training set and tested on the GIM-GLIDE, GIM-DDNM and GIM-VOC test sets to explore the generalization.
+Method
+To address the challenges of generative manipulation, we propose GIMFormer that adopts a dual encoder and decoder architecture.
+Our framework includes several components: the ShadowTracer, the Frequency-Spatial Block (FSB), and the Multi Windowed Anomalous Modeling (MWAM) module.
+Figure 3 gives an overview of the framework.
+For the RGB image input $x$, we first extract its learned trace map $t$.
+Then, both $x$ and $t$ are fed into a two-branch network, where the four-stage structure is used to extract pyramid features $F_{i}$ ($i\in[1,4]$).
+The RGB branch is composed of FSB, Transformer Block (Xie et al.
+2021) and WMAM.
+The tracer branch consists of a Transformer Block and WMAM.
+In the fusion step, the feature rectification module (FRM) and feature fusion module (FFM) (Zhang et al.
+2023) are used for feature fusion.
+The four-stage fused features are forwarded to the decoder for final detection $\hat{y}$ and location $\hat{M}$.
 Figure 3: GIMFormer architecture.
-The ShadowTracer takes the RGB image $x$ to extract the learned manipulation trace map $t$.
-The encoder uses both the RGB input and trace map $t$ to extract pyramidal features $F_{i}$ ($i\in[1,4]$) across four stages.
-The four-stage fused features $F_{i}$ are forwarded to the decoder for generative manipulation detection and location.
-4.1 ShadowTracer
-Prior manipulation detection methods mainly focus on ”cheapfake” scenes and rely on visible traces.
+ShadowTracer extracts trace map $t$ from the input image $x$.
+The encoder combines $x$ and $t$ to generate pyramid features $F_{i}$ across four stages, which are sent to the decoder for manipulation detection and localization.
+ShadowTracer
+Prior manipulation detection methods mainly focus on “cheapfake” and rely on visible traces.
 These artifacts include distortions and sudden changes caused by manipulation of the image structure.
 However, generative tampering makes significant alterations to the content with no apparent frequency or structural inconsistency.
-As shown in Figure 4, these subtle traces are displayed with inherent patterns, not visible traces with inconsistent edges.
+As shown in Figure 4, these subtle traces are displayed with inherent patterns that are not visible traces with inconsistent edges.
 ShadowTracer aims to capture the inherent characteristics and subtle traces of the generative models.
-For a doctored image, our objective is to learn a mapping $g_{\phi}$ to map the tampered image to its latent disturbed pixel values, where $g_{\phi}$ represents a neural network with trainable parameters $\phi$.
+For a manipulated image, our objective is to learn a mapping $g_{\phi}$ to map the tampered image to its latent disturbed pixel values, where $g_{\phi}$ represents a neural network with trainable parameters $\phi$.
 Our key observation is that the differences introduced by generative models in data distribution exhibit inherent patterns, and deep neural networks can attempt to reconstruct these variations.
-At the training stage, we generate pairs of the input image $x_{i}$ and the generative tampered image $G(x_{i})$, the manipulation trace can be calculated by $t_{i}=G(x_{i})-x_{i}$.
-The objective function for training $g_{\phi}$ can be formulated as
-$\min_{\bm{\phi}}\Big{\{}\mathcal{L}_{r}(g_{\bm{\phi}}\left(G\left(\mathbf{x}_{i}\right))\right),t_{i})\Big{\}}$ (1)
+At the training stage, we generate pairs of the image $x_{i}$ and the tampered image $G(x_{i})$, the manipulation trace can be calculated by $t_{i}=G(x_{i})-x_{i}$.
+The objective function for training $g_{\phi}$ can be formulated as :
+$\min_{\boldsymbol{\phi}}\Big{\{}\mathcal{L}_{r}(g_{\boldsymbol{\phi}}\left(G\left(\mathbf{x}_{i}\right))\right),t_{i})\Big{\}}$ (1)
 where $\mathcal{L}_{r}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$.
-Furthermore, the mapping network should satisfy two properties: detecting subtle tampering traces and being robust to various real-world image degradations.
+Furthermore, the mapping network needs to detect subtle tampering traces and be robust to various real-world image degradations.
 For this reason, image pairs are generated by mixing original and manipulated images and incorporating diverse degradation operations at the training stage.
-Specifically, given an input image $I$, we segment a portion and perform generative manipulation to obtain $I_{m}$.
-The alpha blendering [47] is utilized to the original and manipulated images to obscure obvious tampering traces.
-Following this, we subject the images to degradation operations such as blurring, and JPEG compression to obtain the final manipulated image.
-The network is trained on 64$\times$64 pixel patches randomly sampled from the dataset, adapting the loss in eq.
-1.
-4.2 Frequency-Spatial Block
+Specifically, given an input image $I$, we segment the region of interest and perform generative manipulation to obtain $I_{m}$.
+The mix-up (Zhang et al.
+2017a) strategy is utilized to the original and manipulated images to obscure obvious tampering traces.
+Following this, we subject the images to the degradations mentioned above to obtain the final manipulated image.
+The network is trained on 64$\times$64 pixel patches randomly sampled from the dataset and the loss in Eq.
+1 is adopted.
+Frequency-Spatial Block
 When degradation operations are applied, artifacts in manipulated images are tricky to perceive.
-To improve the local expressive ability and effectively harvest discriminative cues in manipulated images, we design a Frequency-Spatial Bloc(FSB) to extract forgery features in the frequency and spatial domains.
-Inspired by the recent work [58, 49, 37, 67], as shown in Figure 3, FSB consists of two branches: a frequency branch and a spatial branch.
+To improve the local expressive and harvest discriminative cues in manipulated images, we design a Frequency-Spatial Block (FSB) to extract forgery features in the frequency and spatial domains.
+Inspired by the recent work (Rao et al.
+2021; Lee-Thorp et al.
+2021; Zhang et al.
+2022), FSB consists of two branches: a frequency branch and a spatial branch as shown in Figure 3.
 In the frequency branch, the input $X$ is converted into the frequency domain $\mathcal{F}_{T}(X)$ using the 2D FFT.
-A learnable filter $G_{i}$ is multiplied to modulate the spectrum and capture the important frequency information.
+A learnable filter $G_{i}$ is multiplied to modulate the spectrum and capture the frequency information.
 Subsequently, the inverse FFT is applied to convert the feature back to the spatial domain, resulting in the extraction of frequency-aware features $X_{\text{f}}$.
-In the spatial branch, the input $X$ is processed through convolutional layers and activated using the LeakyReLU function in a separate network to enhance the expressiveness of the features and obtain refined spatial features $X_{\text{s}}$.
-Then $X_{\text{f}}$ and $X_{\text{s}}$ are concatenated and passed through convolutional layers and the LeakyReLU activation function to obtain enhanced information, which is then combined with the original input $X$ through element-wise summation.
+In the spatial branch, the input $X$ is processed through convolution layers and LeakyReLU function to enhance the expressiveness of the features and obtain refined spatial features $X_{\text{s}}$.
+Then $X_{\text{f}}$ and $X_{\text{s}}$ are concatenated and passed through convolution layers and the LeakyReLU function to obtain enhanced information, which is then combined with the original input $X$ through element-wise summation.
 The total process can be formulated by:
@@
 $X_{out}=\mathrm{Conv_{L}}([X_{\mathrm{f}},X_{\mathrm{s}}])+X,$
-where $\odot$ denotes the Hadamard product , $\mathrm{Conv_{L}}$ denotes convolution with LeakyReLU and $\begin{bmatrix}\cdot\end{bmatrix}$ denotes a concatenation operation.
-Figure 4: Subtle traces are left by generative manipulation devoid of apparent frequency or structural inconsistency.
-It fatally modifies the content and this pixel-level perturbation corrupts the image.
-ShadowTracer captures the intrinsic pattern of the generative tampering and reconstructs the underlying tampering perturbations as priori traces.
-4.3 Multi Windowed Anomalous Modelling Module
+where $\odot$ denotes the Hadamard product , $\mathrm{Conv_{L}}$ denotes convolution with LeakyReLU and $\begin{bmatrix}\cdot\end{bmatrix}$ denotes concatenation.
+Figure 4: Generative manipulation leaves subtle traces, ShadowTracer identifies intrinsic patterns and reconstructs underlying tampering perturbations.
+Multi Windowed Anomalous Modelling Module
 Image manipulation causes discrepancies at the pixel level.
 Genuine pixels are expected to exhibit consistency with neighboring pixels, while manipulated pixels may deviate and display anomalies.
-Former works [62, 35, 28], explore modeling such local inconsistencies.
-To effectively capture the pixel-level inconsistency between the manipulated and real region, we introduce the Multi Windowed Anomalous Modelling module (MWAM) to model these differences at multiple scales for fine-grained features.
-As shown in Figure 3, given the input feature ${F\in H\times W\times C}$, we calculate the difference between each pixel and surrounding pixels within a local window in two branches by Eq.3.
+Motivated by previous works (Wu, AbdAlmageed, and Natarajan 2019; Kong et al.
+2023) that explore local inconsistencies.
+To effectively capture the pixel-level inconsistency between the manipulated and real region, we introduce the Multi Windowed Anomalous Modelling (MWAM) module to model these differences at multiple scales for fine-grained features.
+As shown in Figure 3, for input feature ${F\in H\times W\times C}$, we calculate the difference between pixel and its surroundings within a local window in two branches via Eq.3.
 $D^{k}_{u}[i,j]=(F[i,j]-F^{k}_{u}[i,j])/\sigma^{*},$ (3)
@@
 Then, the obtained $N=3$ different-scale $D^{k}_{a}$ and $D^{k}_{m}$ are concatenated and fed into a convolutional network to obtain an anomaly map $M_{a}$ and $M_{m}$ of the same size as the original input.
-Additionally, the anomaly score mask $\hat{\bm{S_{u}}}\in H\times W$ of the feature is computed using Eq.4.
-$\hat{\bm{f}_{u}}$ $=\mathrm{DW-Conv}\left(\bm{f}\right),$ (4)
-$\hat{\bm{S}_{u}}$ $=\operatorname{Sigmoid}\left(\mathrm{Conv}(C,1)\left(\hat{\bm{f}_{u}}\right)\right),$
-where the $\mathrm{DW}{-\operatorname{Conv}}$ means a $3\times 3$ Depth-Wise convolution layer.
-The element-wise multiplication between anomaly score $\hat{\bm{S}_{u}}$ and the anomaly map $M_{u}$ capture the anomaly information.
+Additionally, the anomaly score mask $\hat{\boldsymbol{S_{u}}}\in H\times W$ of the feature is computed using.
+$\hat{\boldsymbol{f}_{u}}$ $=\mathrm{DConv}\left(\boldsymbol{f}\right),$ (4)
+$\hat{\boldsymbol{S}_{u}}$ $=\operatorname{Sigmoid}\left(\mathrm{Conv}(C,1)\left(\hat{\boldsymbol{f}_{u}}\right)\right),$
+where the $\mathrm{D}{\operatorname{Conv}}$ means a $3\times 3$ Depth-Wise convolution layer.
+The element-wise multiplication between anomaly score $\hat{\boldsymbol{S}_{u}}$ and the anomaly map $M_{u}$ capture the anomaly information.
 Next, we calculate the element-wise summation between the resulting anomaly-aware map and the input feature map $X$ to obtain an anomaly-sensitive feature map.
 The whole process can be described as:
-$\hat{X}=X+\hat{\bm{S}_{a}}\times M_{a}+\hat{\bm{S}_{m}}\times M_{m}$ (5)
-Table 4: Benchmarking IMDL models for manipulation detection and localization for varying mix-generator comprehensive capability.
-The models are trained on the whole GIM training set and test on the correspondence test dataset respectively.
-The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported.
-Params and GFLOPS denote the models’ parameters and computational complexity with Params measured in millions (M).
-$\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
+$\hat{X}=X+\hat{\boldsymbol{S}_{a}}\times M_{a}+\hat{\boldsymbol{S}_{m}}\times M_{m}$ (5)
 Method Params GFLOPS GIM-SD GIM-GLIDE GIM-DDNM
 Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC
-ManTranet[62] 4.0 1009.7 61.08 37.48 80.83 70.99 49.11 83.29 53.99 33.12 74.94
-MVSS-Net[12] 146.9 160.0 56.12 23.17 72.03 61.29 33.12 74.94 49.15 14.11 70.09
-SPAN[25] 15.4 30.9 53.15 35.62 79.28 60.01 39.46 81.21 59.15 32.1 73.55
-PSCC-Net[40] 4.1 107.3 52.28 31.48 83.85 66.52 53.68 86.37 56.33 41.77 85.8
+ManTranet(Wu, AbdAlmageed, and Natarajan 2019) 4.0 1009.7 61.1 37.5 80.8 71.0 49.1 83.3 54.0 33.1 74.9
+MVSS-Net(Dong et al. 2022) 146.9 160.0 56.1 23.2 72.0 61.3 33.1 74.9 49.2 14.1 70.1
+SPAN(Hu et al. 2020) 15.4 30.9 53.2 35.6 79.3 60.0 39.5 81.2 59.2 32.1 73.6
+PSCC-Net(Liu et al. 2022) 4.1 107.3 52.3 31.5 83.8 66.5 53.7 86.4 56.3 41.8 85.8
 […]
 […]
-SegFormer[63] 27.5 41.3 64.25 46.17 83.26 78.11 56.77 88.73 69.29 40.19 84.65
-GIMFormer (Ours) 95.9 96.2 70.92 58.61 88.25 83.89 77.31 95.42 76.72 56.25 88.31
-4.4 Loss Function
-For manipulation detection, we adopt a light-weight backbone in  [57] on the fourth stage feature to calculate the final binary prediction $\hat{y}$.
-For manipulation localization, we utilize the MLP decoder  [63] as the segmentation head to obtain a predicted mask $\hat{M}$.
-Given the ground-truth label y and mask M, we train GIMFormer with the following objective function:
-${\mathcal{L}}={\mathcal{L}}_{cls}(y,{\hat{y}})+\lambda{\mathcal{L}}_{seg}(M,{\hat{M}}),$ (6)
-where both ${\mathcal{L}}_{cls}$ and ${\mathcal{L}}_{seg}$ are binary cross-entropy loss, and $\lambda=1$ is a balancing hyperparameter.
-4.5 Implementation Details.
-Our approach includes two separate training steps.
-First, we train the ShadowTracer using a synthesized dataset of ImageNet.
-This training process follows a similar data generation method as described in Section 4.1.
-Then, we train the encoder and decoder of the model according to the two settings in GIM, as described in Section 3.4.
-We train our models on eight V100 GPUs with an initial learning rate (LR) of $6e^{-5}$ which is scheduled by the poly strategy with power 0.9 over 40 epochs.
-The optimizer is AdamW  [41] with epsilon $1e^{-8}$ weight decay $1e^{-2}$, and the batch size is 4 on each GPU.
-Table 5: Benchmarking IMDL models for manipulation detection and localization for varying cross-generator generalization capability.
-The models are trained on the GIM-SD training set and tested on all the test sets to explore the generalization ability.
-The detection Cls.Acc(%) and localization AUC(%) and F1(%) are reported.
-$\ddagger$ indicates that the original paper did not provide the code, we reproduced the corresponding code and evaluated it under the same settings
-Method GIM-SD GIM-SD(VOC) GIM-GLIDE GIM-DDNM
+IML-VIT(Ma et al. 2024) 91.0 136.0 65.2 53.9 84.3 71.3 69.3 92.1 71.2 53.9 85.7
+SegFormer(Xie et al. 2021) 27.5 41.3 64.3 46.2 83.3 78.1 56.8 88.7 69.3 40.2 84.6
+GIMFormer (Ours) 95.9 96.2 70.9 58.6 88.2 83.9 77.3 95.42 76.7 56.3 88.3
+Table 4: Benchmarking IMDL models to evaluate performance. Cls.Acc(%), AUC(%) and F1(%) Params(M) are reported.
+Method GIM-SD GIM-VOC GIM-GLIDE GIM-DDNM
 Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC Cls.Acc F1 AUC
-ManTranet[62] 73.12 43.18 80.18 63.18 27.37 72.74 58.21 24.53 74.61 39.52 16.8 58.31
-MVSS-Net[12] 56.12 25.13 82.15 56.30 21.2 73.17 53.12 23.17 73.11 48.12 10.11 50.15
-SPAN[25] 52.15 47.32 56.93 52.17 29.86 55.17 50.03 30.14 56.08 44.17 14.30 56.67
-PSCC-Net[40] 58.24 48.36 87.29 54.79 39.36 83.80 51.11 29.60 79.76 40.52 4.40 48.52
+ManTranet(Wu, AbdAlmageed, and Natarajan 2019) 73.1 43.2 80.2 63.2 27.4 72.7 58.2 24.5 74.6 39.5 16.8 58.3
+MVSS-Net(Dong et al. 2022) 56.1 25.1 82.2 56.3 21.2 73.2 53.2 23.17 73.1 48.1 10.1 50.2
+SPAN(Hu et al. 2020) 52.2 47.3 56.9 52.2 29.9 55.2 50.0 30.1 56.1 44.2 14.3 56.7
+PSCC-Net(Liu et al. 2022) 58.2 48.4 87.3 54.8 39.4 83.8 51.1 29.6 79.8 40.5 4.4 48.5
 […]
 […]
-SegFormer[63] 71.93 53.08 87.11 60.03 28.17 80.97 54.11 21.03 75.12 50.21 4.70 51.33
-GIMFormer (Ours) 78.96 61.75 90.61 67.19 50.01 84.01 67.01 39.27 81.02 50.23 19.37 60.12
-4.6 Comparison with state-of-the-art methods
-We compare our methods with various state-of-the-art IMDL methods  [62, 12, 25, 40, 19, 58] on generative manipulation detection and location.
-In addition, the vanilla SegFormer (MiT-B2) [63] is also compared, since our method is based on that architecture.
-As shown in Table 4 and Table 5, these methods are tested on various settings to verify the performance and generalization.
+IML-VIT(Ma et al. 2024) 77.1 58.1 88.9 54.3 47.2 83.2 52.1 23.1 71.1 50.3 7.2 52.3
+SegFormer(Xie et al. 2021) 71.9 53.1 87.1 60.0 28.2 81.0 54.1 21.0 75.1 50.2 4.7 51.3
+GIMFormer (Ours) 78.9 61.7 90.6 67.2 50.0 84.0 67.0 39.3 81.0 52.3 19.4 60.1
+Table 5: Benchmarking IMDL models to evaluate generalization. Cls.Acc(%), AUC(%) and F1(%) are reported.
+Loss Function
+For detection, we adopt a lightweight backbone in  (Wang et al.
+2020) on the fourth stage feature for binary prediction $\hat{y}$.
+For localization, we utilize the MLP decoder  (Xie et al.
+2021) as the segmentation head to obtain a predicted mask $\hat{M}$.
+Given the ground-truth label $y$ and mask $M$, we train GIMFormer with the following objective function:
+${\mathcal{L}}={\mathcal{L}}_{cls}(y,{\hat{y}})+{\mathcal{L}}_{seg}(M,{\hat{M}}),$ (6)
+where both ${\mathcal{L}}_{cls}$ and ${\mathcal{L}}_{seg}$ are binary cross-entropy loss.
+Implementation Details
+Our approach includes two separate training steps.
+First, we train the ShadowTracer using the dataset generated from ImageNet.
+This training process follows a similar data generation method as mentioned in the former chapter.
+Then, we train the encoder and decoder of the model according to the two settings in GIM, as described in the previous section.
+We train our models on 8 V100 GPUs with an initial learning rate of $6e^{-5}$ which is scheduled by the poly strategy with power 0.9 over 20 epochs.
+The optimizer is AdamW  (Loshchilov and Hutter 2017) with epsilon $1e^{-8}$ weight decay $1e^{-2}$, and the batch size is 4 on each GPU.
+Comparison with state-of-the-art methods
+We compare our methods with various state-of-the-art IMDL methods 111 $\ddagger$ indicates that the original paper does not provide the code, we reproduce the code and evaluate it under the same settings.
+and the vanilla SegFormer (MiT-B2) (Xie et al.
+2021).
+As shown in Table 4 and Table 5, these methods are tested in two settings to evaluate the performance.
 Note that some methods are not explicitly designed for image-level detection, in which case we use the maximum of the prediction map as the detection statistic.
 All methods are immersed in the same implementation details.
-Cross-generator generalization capability comparison.
-Table 5 reports the generalization performance of all the methods mentioned.
-The results show that GIMFormer outperforms all other methods by a significant margin in both in-domain detection ability and cross-domain generalization ability.
-For in-domain detection ability, our method efficiently catches the subtle artifacts inherent in generative and accurately localizes them.
-Meanwhile, other methods may encounter confusion as they attempt to learn specific content, potentially leading to challenges in accurately detecting and localizing generative tampering patterns.
-For cross-domain generalization ability, GIMFormer achieves state-of-the-art performance.
-It demonstrates cross-detection capability in detecting manipulation using different generative models, as shown in the results of the GIM-GLIDE and GIM-DDNM testsets.
-Besides GIMFormer works well on Data generated by the same generator from different distributions, while existing methods are confused and have an obvious performance drop, as shown in the results of the GIM-SD(VOC) testset.
+More experiments are included in the Appendix.
+Mix-generator comparison.
+Table 4 reports the performance of all methods.
+GIMFormer outperforms other methods, demonstrating its superior ability to identify generative manipulation.
+Existing methods exhibit low pixel-level F1 scores on this benchmark, indicating that tampering areas are not accurately identified.
 The qualitative results for visual comparisons are illustrated in Figure 5.
-Mix-generator comprehensive capability comparison Table 5 reports the comprehensive performance of all the methods mentioned.
-GIMFormer also outperforms all other methods on all the test sets, demonstrating its superior ability to identify generative tampering and its comprehensive performance.
-Existing methods have a low pixel-level F1 score on this benchmark, which means that tampering areas cannot be accurately identified.
+Cross-generator comparison.
+Table 5 reports the generalization of all the methods.
+The results show that GIMFormer outperforms all other methods in both in-domain and cross-domain.
+For in-domain experiments, our method catches the subtle artifacts inherent in manipulation and accurately localizes them.
+Other methods may encounter confusion as they attempt to learn specific content, potentially leading to challenges in accurately detecting and localizing generative tampering patterns.
+For cross-domain experiments, GIMFormer demonstrates well generalization in detecting manipulation using different generative models, as shown in the results of the GIM-GLIDE and GIM-DDNM testsets.
+Besides GIMFormer works well on data generated by the same generator from different distributions, while existing methods have an obvious performance drop, as shown in the results of the GIM-VOC testset.
 The qualitative results for visual comparisons are illustrated in Figure 6.
-Figure 5: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods comprehensive capability.
-From top to bottom, we show examples from the authentic image, GIM-SD, GIM-GLIDE, GIM-DDNM
-Figure 6: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods generalization capability. From top to bottom, we show examples from GIM-SD, GIM-GLIDE, GIM-DDNM
-4.7 Ablation Analysis
-Effectiveness of proposed module.
-We consider a simple baseline proposed in  [63] and gradually integrate new key components.
+Figure 5: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods performance.
+Figure 6: Qualitative results on GIM of comparing GIMFormer with state-of-the-art methods generalization capability.
+Ablation Analysis
+Effectiveness of the proposed module.
+We consider a simple baseline proposed in  (Xie et al.
+2021) and gradually integrate new key components.
 Experiments are carried out on the GIM-GLIDE testset.
@@
 The results verify that ShadowTracer, FSB and MWAM effectively improve the performance of the baseline model.
-Table 6: Ablation results on GIM-GLIDE test dataset using different variants of GIMFormer. All detection Cls.Acc(%), localization AUC(%), and F1(%) scores are reported.
 Variants Cls.Acc F1 AUC
-Baseline 78.11 56.77 88.73
-+ST 79.96 67.12 91.99
-+ST+WMAM 80.19 73.41 94.01
-+ST+WMAM+FSB 83.89 77.31 95.42
-Table 7: Generalization Experiments of ShadowTracer in percentage (%). ST denotes ShadowTracer and ST(SD) denotes ShadowTracer trained on data generated by Stable-Diffusion.
+Baseline 78.1 56.8 88.7
++ShadowTracer 79.9 67.1 92.0
++ShadowTracer+WMAM 80.2 73.4 94.0
++ShadowTracer+WMAM+FSB 83.9 77.3 95.4
+Table 6: Ablation results of GIMFormer variants. Cls.Acc(%), AUC(%) and F1(%) scores are reported.
 Variants GIM-GLIDE GIM-DDNM
 Cls.Acc F1 AUC Cls.Acc F1 AUC
-GIMFormer w ST(SD) 83.1 78.1 95.4 77.4 58.8 89.4
-GIMFormer w/o ST 80.0 68.9 91.3 73.1 51.3 86.1
+GIMFormer w ShadowTracer 83.1 78.1 95.4 77.4 58.8 89.4
+GIMFormer w/o ShadowTracer 80.0 68.9 91.3 73.1 51.3 86.1
+Table 7: Generalization Experiments of ShadowTracer. Cls.Acc(%), AUC(%) and F1(%) scores are reported.
 Generalization of ShadowTracer.
-We initiate the training process of ShadowTracer using data generated by Stable Diffusion and subsequently hold its weights fixed.
-Following this pretraining phase, we proceed to train the backbone on both the GIM-GLIDE and GIM-DDNM trainsets, with and without the incorporation of ShadowTracer.
-The detection result Cls.Acc and location results in F1 and AUC are presented in Table 7, revealing that leveraging the pretrained ShadowTracer significantly enhances performance in cross-generator IMDL tasks.
-Notably, in the GIM-GLIDE dataset with the same manipulation type, leveraging ShadowTracer leads to a remarkable 9% enhancement in F1 and a 4% increase in AUC for location results.
-Additionally, the benefits extend to the detection aspect, manifesting as a 3% increase in accuracy.
-In the GIM-DDNM dataset with various manipulation types, there is a 7% enhancement in F1, a 3% increase in AUC for location results and a 4 % increase in accuracy.
-5 Conclusion
-We address the challenge of detecting and locating generative-based manipulation and provide a reliable database (GIM) for AIGC security.
-The proposed dataset leverages multiple mainstream generators and tampering methods to provide a variety of generative manipulation data.
-Additionally, We introduce GIMFormer, a novel transformer-based IMDL framework.
-ShadowTracer is designed to catch subtle artifacts in generative tampering.
-While the Frequency-Spatial Block gathers frequency and spatial information, the Multi Windowed Anomalous Modelling module captures pixel-level inconsistencies at multiple scales for fine-grained features.
-Extensive experiments demonstrate the superior performance of our model, which achieves SoTA results.
-Limations.
-The GIM dataset shares classes with ImageNet and VOC, but it may not encompass future emerging objects due to the evolving variety in the real world.
-Fine-tuning pre-trained models on new object data can address this issue.
-Current research primarily centers on image manipulation.
-The emergence of realistic video generation models like SORA [5] presents fresh challenges as AI manipulation extends into videos.
-Future plans involve expanding research to address the complexities of video manipulation.
-Supplementary Document
-In this supplementary document, we include many details of our work: (1) The details of the proposed GIM Datset and compare with related work (Sec.
-6) (2) The details of Architecture and implementation (Sec.
-7).
-(3) The additional ablation analysis of GIMFormer(Sec.
-8) (4) More Visualizations on GIM dataset (Sec.
-9).
-(5)The societal impact of our work (Sec.
-10).
-(6) The ethics statement of our work (Sec.11).
-6 GIM Dataset
-6.1 GIM Dataset Configuration
+To evaluate the generalization of ShadowTracer, we train ShadowTracer using data generated by Stable Diffusion and subsequently hold its weights fixed.
+We proceed to train the backbone on the GIM-GLIDE and GIM-DDNM trainsets, with and without the incorporation of ShadowTracer.
+The result in Table 7 reveals that leveraging the pretrained ShadowTracer increases performance in cross-generator IMDL tasks.
+Across GIM-GLIDE and GIM-DDNM, ShadowTracer achieves up to 9% F1, 4% AUC and 4% accuracy improvements.
+Conclusion
+We address the challenge of detecting and locating generative manipulation and provide a reliable database GIM for AIGC security.
+This dataset leverages multiple generators to provide diverse generative manipulation data.
+Based on this, we design a benchmark for IMDL methods with two settings.
+We also introduce GIMFormer, a novel transformer-based IMDL framework.
+Extensive experiments demonstrate that GIMFormer achieves SOTA performance.
+Appendix
+In this Appendix, we include more details of our work: (1) The details of the proposed GIM Datset and compare with related work (2) Details of the GIMFormer architecture and implementation (3) The additional ablation analysis of GIMFormer (4) More Visualizations on GIM dataset.
+(5)The societal impact of our work .
+(6) The ethics statement of our work.
+Appendix A GIM Dataset
+GIM Dataset Configuration
 The specific quantities and details of the dataset are presented in Table 9, GIM has a total number of 1,140k generative manipulated images with their corresponding origin images.
 Moreover, the dataset can be split into subsets.
-GIM-SD-all utilizes the training set from Imagenet-1k [10] and employs the Stable Diffusion [53] generator for image manipulation dedicated to manipulation research.
+GIM-SD-all utilizes the training set from Imagenet-1k (Deng et al.
+2009) and employs the Stable Diffusion generator for image manipulation dedicated to manipulation research.
 Through the analysis of the data scale, GIM-SD, GIM-GLIDE and GIM-DDNM individually select a random set of 100 classes from the Imagenet-1k training dataset and the entire test data to generate their respective training and test sets to benchmark the image manipulation detection and location (IMDL)methods.
-GIM-SD (VOC) is a cross-data-distribution test set constructed using the test set from the PASCAL VOC [14] dataset and Stable Diffusion techniques.
+GIM-VOC is a cross-data-distribution test set constructed using the test set from the PASCAL VOC (Everingham et al.
+2010) dataset and Stable Diffusion techniques.
 Table 8: Parameters for GIM degradation. Three degradation methods are employed with variable parameters to emulate real-world scenarios.
@@
 GIM-DDNM removal 125K 50K
-GIM-SD(VOC) reprint - 4.5K
+GIM-VOC reprint - 4.5K
 To simulate real-world scenarios and comprehensively evaluate the IMDL methods, GIM incorporates three degradation methods (JPEG compression, Gaussian blur, and downsampling).
@@
 The GIM data format consists of JPG images paired with PNG labels.
-Filenames ending with "_f" indicate tampered images.
-The masks segmented by SAM [33] serve as ground truth for manipulation detection, containing two labels denoted by 0 and 255, representing the original and manipulated categories, respectively.
-6.2 Dataset Comparison.
+Filenames ending with ”_f” indicate tampered images.
+The masks segmented by SAM (Kirillov et al.
+2023b) serve as ground truth for manipulation detection, containing two labels denoted by 0 and 255, representing the original and manipulated categories, respectively.
+Dataset Comparison.
 GIM focuses on generative local manipulation in natural images and provides manipulated region labels, specifically for image manipulation detection and localization tasks.
@@
 These datasets commonly employ manual or random manipulation to generate data.
-AutoSplice [29], HiFi-IFDL[20], CoCoGlide  [19] dataset utilizes generative algorithms for image manipulation.
+AutoSplice (Jia et al.
+2023), HiFi-IFDL(Guo et al.
+2023), CoCoGlide  (Guillaro et al.
+2023) dataset utilizes generative algorithms for image manipulation.
 However, they rely on one type of manipulation and suffer from a small data volume.
@@
 # Authetic Image # Forged Image Splicing Copy-move Removal Removal Generation
-Columbia Gray [43] 933 912
-Columbia Color [24] 183 180
-MICC-F2000 [1] 1,300 700
-VIPP Synth.  [1] 4,800 4,800
-CASIA V1.0  [13] 800 921
-CASIA V2.0  [13] 7,200 5,123
-Wild Web [66] 90 9,657
-NC2016  [17] 560 564
-NC2017  [17] 2,667 1,410
-MFC2018  [17] 14,156 3,265
-MFC2019  [17] 10,279 5,750
-PS-Battles [22] 11,142 102,028
-DEFACTO  [42] 0 229,000
-IMD2020 [45] 35,000 35,000
-SP COCO  [36] 0 200,000
-CM COCO  [36] 0 200,000
-CM RAISE  [36] 0 200,000
-HiFi-IFDL[20] - 1,000,00*
-CoCoGlide [19] 0 512
-AutoSplice  [29] 3,621 2,273
+Columbia Gray (Ng, Chang, and Sun 2004) 933 912
+Columbia Color (Hsu and Chang 2006) 183 180
+MICC-F2000 (Amerini et al. 2011) 1,300 700
+VIPP Synth.  (Amerini et al. 2011) 4,800 4,800
+CASIA V1.0  (Dong, Wang, and Tan 2013) 800 921
+CASIA V2.0  (Dong, Wang, and Tan 2013) 7,200 5,123
+Wild Web (Zampoglou, Papadopoulos, and Kompatsiaris 2015) 90 9,657
+NC2016  (Guan et al. 2019) 560 564
+NC2017  (Guan et al. 2019) 2,667 1,410
+MFC2018  (Guan et al. 2019) 14,156 3,265
+MFC2019  (Guan et al. 2019) 10,279 5,750
+PS-Battles (Heller, Rossetto, and Schuldt 2018) 11,142 102,028
+DEFACTO  (Mahfoudi et al. 2019) 0 229,000
+IMD2020 (Novozamsky, Mahdian, and Saic 2020) 35,000 35,000
+SP COCO  (Kwon et al. 2022) 0 200,000
+CM COCO  (Kwon et al. 2022) 0 200,000
+CM RAISE  (Kwon et al. 2022) 0 200,000
+HiFi-IFDL(Guo et al. 2023) - 1,000,00*
+CoCoGlide (Guillaro et al. 2023) 0 512
+AutoSplice  (Jia et al. 2023) 3,621 2,273
 GIM 1,140,000 1,140,000
-7 Implementation Details
-7.1 Architecture
+Appendix B Implementation Details
+Architecture
 […]
-7.2 Training Process
+Training Process
 We conduct our experiments with PyTorch 1.7.0. All models are trained on a node with 8 or 4 V100 GPUs.
 For the training of the ShadowTracer network $g_{\phi}$, We randomly selected 20,000 authentic images from ImageNet and correspondingly generated manipulated images.
-During comprehensive performance verification, all three generators (Stable Diffusion [52], GLIDE [44], DDNM [59])are used, whereas for generalization verification, only Stable Diffusion is employed.
+During comprehensive performance verification, all three generators (Stable Diffusion (Rombach et al.
+2022), GLIDE (Nichol et al.
+2021), DDNM (Wang, Yu, and Zhang 2022))are used, whereas for generalization verification, only Stable Diffusion is employed.
 Throughout the training, the alpha blending parameter ranges between 0.5 and 1.0, randomly producing blended images subjected to three degradation types (JPEG compression, downsampling, Gaussian blur).
@@
 Training is conducted for roughly 300,000 iterations with a batch size of 64.
-An Adam [31] optimizer is employed, initialized with a learning rate of 0.001.
+An Adam (Kingma and Ba 2014) optimizer is employed, initialized with a learning rate of 0.001.
 For the training of the GIMFormer backbone network.
@@
 The images are augmented by random resize with a ratio of 0.5–2.0, random horizontal flipping.
-8 Additional Ablation Analysis
+Appendix C Additional Ablation Analysis
 The robustness of GIMFormer.
@@
 {$D_{a\&k}^{11},D_{a\&k}^{21},D_{a\&k}^{31},D_{a\&k}^{41}$} 78.11 61.15 90.71
-9 More Visualizations on GIM
+Appendix D More Visualizations on GIM
 In this section, we present additional visualizations of the GIM dataset in Figure  7,  8, and  9, showcasing the data visualization from GIM-SD, GIM-GLIDE, and GIM-DDNM respectively.
-10 Societal Impact
-Our research yields a positive societal impact on the community by focusing on addressing the challenge of detecting and locating generative-based manipulations.
-We introduce a reliable database, GIM, aimed at enhancing the security of AI-generated content (AIGC).
-This database facilitates the training of Image Manipulation Detection and Localization (IMDL)methods on GIM, allowing them to generalize across various scenarios.
-Consequently, our algorithm GIMFormer fosters increased trust among the general public in our society regarding media content.
-GIM and GIMFormer are beneficial for digital media forensics, especially generative manipulation with real-world degradations.
-11 Ethics Statement
-Our GIM is based on ImageNet and VOC.
-No additional personally identifiable information or sensitive personally identifiable information is introduced during the production of fake images in the GIM dataset.
-During the dataset production, we do not introduce extra information containing exacerbated bias against people of a certain gender, race, sexuality, or who have other protected characteristics.
-The ethical issues in the ImageNet and VOC datasets have been discussed in previous works.
-Crawford et al.
-[8] discuss issues with ImageNet.
-The first issue is the political nature of all taxonomies or classification systems, where terms like "male" and "female" are considered "natural," while "hermaphrodite" is offensively placed within the branch of Person > Sensualist > Bisexual alongside "pseudohermaphrodite" and "switch hitter" categories.
-The second issue concerns offensive images of real people, while the third is the use of people’s photos without their consent by ImageNet creators.
+These figures illustrate the diverse data representations within the GIM dataset, showcasing the unique characteristics brought by different generative models and manipulation techniques.
 Figure 7: Additional visualizations of GIM-SD.
@@
 From left to right, the sequence is as follows: original image, tampered area, generative manipulated image, original image, tampered area, generative manipulated image.
+Appendix E Societal Impact
+Our research contributes positively to society by addressing the challenge of detecting and locating generative-based manipulations, ultimately enhancing the security of AI-generated content (AIGC).
+The introduction of the reliable GIM database aims to bolster the effectiveness of Image Manipulation Detection and Localization (IMDL) methods, allowing for generalization across diverse scenarios.
+As a result, our algorithm GIMFormer instills greater trust among the general public in our society regarding media content authenticity.
+The utilization of GIM and GIMFormer proves advantageous for digital media forensics, particularly in addressing generative manipulation with real-world degradations.
+The practical applications of the GIM dataset have the potential to create a significant social impact.
+For instance, the training of detectors on GIM enables users to identify AI-generated content, leading to consequential actions such as identifying academic dishonesty.
+Widely disseminating and utilizing the GIM dataset for IMDL algorithms training could pave the way for its integration into specialized software designed for this purpose, thereby enhancing the accuracy and reliability of algorithms.
+However, it is essential to note that models trained on this dataset may exhibit undesirable tendencies towards novel objects in the real world, necessitating careful consideration for practical implementation despite their demonstrated generalizability.
+Appendix F Ethics Statement
+Our GIM is based on ImageNet and VOC.
+No additional personally identifiable information or sensitive personally identifiable information is introduced during the production of fake images in the GIM dataset.
+During the dataset production, we do not introduce extra information containing exacerbated bias against people of a certain gender, race, sexuality, or who have other protected characteristics.
+The ethical issues in the ImageNet and VOC datasets have been discussed in previous works.
+Crawford et al.
+(Crawford and Paglen 2021) discuss issues with ImageNet.
+The first issue is the political nature of all taxonomies or classification systems, where terms like ”male” and ”female” are considered ”natural,” while ”hermaphrodite” is offensively placed within the branch of Person ¿ Sensualist ¿ Bisexual alongside ”pseudohermaphrodite” and ”switch hitter” categories.
+The second issue concerns offensive images of real people, while the third is the use of people’s photos without their consent by ImageNet creators.
+Appendix G Limitations
+The GIM dataset shares classes with ImageNet and VOC, but it may not encompass future emerging objects due to the evolving variety in the real world.
+Fine-tuning pre-trained models on new object data can address this issue.
+Current research primarily centers on image manipulation.
+The emergence of realistic video generation models like SORA presents fresh challenges as AI manipulation extends into videos.
+Future plans involve expanding research to address the complexities of video manipulation.
+Appendix H Acknowledgments
+This work is partially supported by NSFC ( No. 62402318, 62376153, 24Z990200676)